### v1

# Advances in Nonmonotone Proximal Gradient Methods merely with Local Lipschitz Assumptions in the Presense of Kurdyka–Łojasiewicz Property: A Study of Average and Max Line Search

Abstract. The proximal gradient method is a standard approach to solve the composite minimization problems where the objective function is the sum of a continuously differentiable function and a lower semicontinuous, extended-valued function. For both monotone and nonmonotone proximal gradient methods, the convergence theory has traditionally replied heavily on the assumption of global Lipschitz continuity. Recent works have shown that the monotone proximal gradient method, even when the local Lipschitz continuity (rather than global) is assumed, converges to the stationarity globally in the presence of Kurdyka–Łojasiewicz Property. However, how to extend these results from monotone proximal gradient method to nonmonotone proximal gradient method (NPG) remains an open question. In this manuscript, we consider two types of NPG: those combined with average line search and max line search, respectively. By partitioning of indices into two subsets, one of them aims to achieve a decrease in the functional sequence, we establish the global convergence and rate-of-convergence (same as the monotone version) results under the KL property, merely requiring the local Lipschitz assumption, and without an a priori knowledge of the iterative sequence being bounded. When our work is almost done, we noticed that [17] presented the analogous results for the NPG with average line search, whose partitioning of index set is totally different with ours. Drawing upon the findings in this manuscript and [17], we confidently conclude that the convergence theory of NPG is independent on the specific partitioning of the index set.
Keywords. Composite problems $\cdot$ Nonmonotone proximal gradient methods $\cdot$ Kurdyka-Łojasiewicz property $\cdot$ Local Lipschitz continuity
AMS subject classifications. 49J52, 90C26, 90C30
1 Introduction
Let us consider
$\min_{x}\ q(x):=f(x)+g(x)\qquad\text{s.t.}\quad x\in\mathbb{X},$ (Q)
where $\mathbb{X}$ is an Euclidean space, i.e., real and finite-dimensional Hilbert space, $f:\mathbb{X}\to\mathbb{R}$ is continuously differentiable, $g:\mathbb{X}\to\overline{\mathbb{R}}$ is assumed to be merely lower semicontinuous. We claim that the unconstrained composite optimization problem (Q) is in a totally nonconvex setting. This type of optimization problems have a wide range of applications in practice, like machine learning, image processing, and data science [4, 11, 8].
In this manuscript, we are interested in two classes of nonmonotone proximal gradient method to solve (Q), where the nonmonotone line search technique based on monitoring the weighted average of the objective value at all iterates by Zhang and Hager [28], and the maximum objective value attained from the latest iterates by Grippo et al. [13], respectively. More precisely, some iterate $x^{k+1}\in\operatorname{dom}q$ is acceptable provided that the corresponding objective $q(x^{k+1})$ is less than the merit
$\Phi_{k}:=(1-p)\Phi_{k-1}+pq(x^{k})\qquad\text{and}\qquad q(x^{l(k)}):=\max_{j=0,\ldots,\min\{m,k\}}q(x^{k-j})$
for the average and max taste, respectively. Note that $m\in\mathbb{N}$ and $p\in(0,1]$ (usually) are constants in order to control the level of nonmonotonicity, i.e., for the smaller $p$, or the larger $m$, the corresponding metric function can attain larger values, hence imposing a weaker condition for acceptance of $x^{k+1}$ compared to the monotone proximal gradient methods [12]. The values of $p=1$ and $m=0$ lead to the monotone behavior, i.e., the nonmotone proximal gradient method degenerates into the monotone one, please see [18, Algorithm 3.1] and [15, Algorithm 3.1].
The convergence theory associated with (monotone and nonmonotone) proximal gradient methods mostly require the derivative of the smooth part of the objective function to be global Lipschitz continuous, which, however, is quite restrictive assumption in practical scenarios. [15, Examples 3.6 and 3.7], as two classical examples, revealed that the global Lipschitz assumption on the gradient of $f$ is typically violated, whereas the local Lipschitz condition is often satisfied. [18] readdressed the classical monotone proximal gradient method [18, Algorithm 3.1] and NPG with max line search [18, Algorithm 4.1], and provided (subsequential) convergence results only with local Lipschitz assumption (not global any more). As for the convergence of the whole sequence generated by the proximal gradient method with local Lipschitz assumption, [15] is the first one to reveal this mystery for monotone proximal gradient in the presence of the Kurdyka-Łojasiewicz property of the objective function. [12] considered an adaptive nonmonotone proximal gradient scheme based on an averaged metric function and established (subsequential) convergence results with local Lipschitz assumption of the gradient of $f$, where the global worst-case rates for the iterates and a (subsequential) stationarity measure were also derived. Note that the numerical behaviours and competitiveness has already been investigated in many literature before, for example see [26, 12, 25, 20, 21, 27], this manuscript focuses more on theoretical aspects of the method (without numerical experiments), showing qualitative and even quantitative properties of the NPG in a more general framework which allows for applications e.g. in the context of multiplier-penalty methods [16]. More precisely, we, in this manuscript, under the same assumptions, consider the convergence theory of the entire sequence generated by NPG with both average and max backtracking line search when KL property being employed.
Note that the corresponding convergence results about NPG with max line search has been analysed in [25], where an auxiliary sequence is necessarily prerequisite to be convergent, which, however, is not needed at all in our technique, moreover, where the boundedness of subdifferential of the objective function is needed for all iterates, which, to the best of our knowledge, fails to be satisfied under the local Lipschitz continuous assumption. However, in this manuscript, no any prerequisite is assumed and the corresponding convergence and rate-of-convergence results obtained in [25] can be also achieved merely by requiring the local Lipschitz continuity of $\nabla f$.
The primary challenge posed by NPG, in fact, is that the generated functional sequence may not be decreasing, to date, whereas the (sufficient) descent property seems to be necessary for the technical proof when using KL property [2, 3, 7, 8]. To the best of our knowledge, nearby all existing studies addressing this issue reply on the partition of the index sets [25, 17, 26]. Specifically, these works divided the indices into two subsets: one subset ensures a decrease in the functional sequence, while the complement accounts for the remaining cases. The recent paper [17] considered the convergence results of the NPG with average line search under the KL property, merely requiring the local Lipschitz continuity of $\nabla f$, where the index sets are defined as
$K_{1}:=\{k\in\mathbb{N}\,|\,q(x^{k})\leq\Phi_{k+m}\},\quad K_{2}:=\{k\in\mathbb{N}\,|\,q(x^{k})>\Phi_{k+m}\},$
where $m$ is some predetermined integer relying on $p_{\min}$ to facilitate the convergence analysis. This partitioning originates from [26] and the subsequent convergence analysis heavily depend on the techniques therein, where the key component is exploiting the quantitative relationship between $\sum_{i=k}^{k+m-1}\sqrt{\Phi_{i-1}-\Phi_{i}}$ and $\sqrt{\Phi_{i-1}-\Phi_{i}}$, see [17, Lemma 4.4] and [26, Theorem 2.1]. To be honest, we and [17] happened to consider the same topic almost simultaneously. In this manuscript, for the NPG with average line search, to the best of our knowledge, we are the first one to propose a different partitioning of the index sets from [26]:
$S:=\{k\in\mathbb{N}\,|\,q(x^{k})-\Phi_{k+1}\leq\frac{\mu}{2}\|x^{k+1}-x^{k}\|^{2}\},\quad\overline{S}:=\mathbb{N}\setminus S,$
where $\mu\in(0,\frac{1}{2}\delta{p_{\min}\gamma_{\min}}]$ is a constant, and $p_{\min}$ and $\gamma_{\min}$ are parameters derived from the algorithm. Actually, different partitioning of the index sets necessitates the different techniques for proof. However, our approach yields more natural and concise proofs compared to [17], as it eliminates the need for the mentioned-above quantitative relationship ([17, Lemma 4.4] and [26, Theorem 2.1]). Our analysis reproduces the corresponding convergence and rate-of-convergence presented in [17]. Furthermore, [17] leaves the theoretical analysis of NPG methods with max line search for future work, we provide a comprehensive analysis in this manuscript. The decision to present this part of work (NPG with average line search), despite its similarities to [17], is also justified by a noteworthy finding: both works demonstrate that the convergence results of nonmonotone proximal gradient methods are independent on the specific partitioning of the index sets. This insight contributes a significant advancement to the understanding of this method.
This manuscript is organized as follows: We first recall some background knowledge in Section 2. We then present the basic properties and convergence results of the NPG with average line search and with max line search in Section 3 and Section 4, respectively. Section 5 concludes this manuscript.
2 Preliminaries
[…]
$\operatorname{dist}(x,A):=\inf\{\left\|y-x\right\|\,|\,y\in A\}$
denotes the distance of the point $x$ to the set $A$ with $\operatorname{dist}(x,\emptyset):=\infty$. For given $\varepsilon>0$, $B_{\varepsilon}(x):=\{y\in\mathbb{X}\,|\,\left\|y-x\right\|\leq\varepsilon\}$ denotes the closed $\varepsilon$-ball around $x$.
The continuous linear operator $f^{\prime}(x)\colon\mathbb{X}\to\mathbb{R}$ denotes the derivative of the continuously differentiable function $f\colon\mathbb{X}\to\mathbb{R}$ at $x\in\mathbb{X}$, and we will make use of $\nabla f(x):=f^{\prime}(x)^{*}1$ where $f^{\prime}(x)^{*}\colon\mathbb{R}\to\mathbb{X}$ is the adjoint of $f^{\prime}(x)$. This way, $\nabla f$ is a mapping from $\mathbb{X}$ to $\mathbb{X}$.
We further say that a sequence $\{x^{k}\}\subset\mathbb{X}$ converges Q-linearly to $x^{*}\in\mathbb{X}$ if there is a constant $c\in(0,1)$ such that the inequality
$\|x^{k+1}-x^{*}\|\leq c\|x^{k}-x^{*}\|$
holds for all sufficiently large $k\in\mathbb{N}$. Furthermore, $\{x^{k}\}$ is said to converge R-linearly to $x^{*}$ if we have
$\limsup_{k\to\infty}\|x^{k}-x^{*}\|^{1/k}<1.$
Note that this R-linear convergence holds if there exist constants $\omega>0$ and $\mu\in(0,1)$ such that $\|x^{k}-x^{*}\|\leq\omega\mu^{k}$ holds for all sufficiently large $k\in\mathbb{N}$, i.e., if the expression $\|x^{k}-x^{*}\|$ is dominated by a Q-linearly convergent null sequence.
Let us fix a merely lower semicontinuous function $q\colon\mathbb{X}\to\overline{\mathbb{R}}$ and pick $x\in\operatorname{dom}q$ where $\operatorname{dom}q:=\{x\in\mathbb{X}\,|\,q(x)<\infty\}$ denotes the domain of $q$. Then the set
^$\widehat{\partial}q(x):=\left\{\eta\in\mathbb{X}\,\middle|\,\liminf\limits_{y\to x,\,y\neq x}\frac{q(y)-q(x)-\langle\eta,y-x\rangle}{\left\|y-x\right\|}\geq 0\right\}$
is called the regular (or Fréchet) subdifferential of $q$ at $x$. Furthermore, the set
$\partial q(x):=\left\{\eta\in\mathbb{X}\,\middle|\,\begin{aligned} &\exists\{x^{k}\},\{\eta^{k}\}\subset\mathbb{X}\colon\\ &\qquad x^{k}\to x,\,q(x^{k})\to q(x),\,\eta^{k}\to\eta,\,\eta^{k}\in\widehat{\partial}q(x^{k})\,\forall k\in\mathbb{N}\end{aligned}\right\}$
is well known as the limiting (or Mordukhovich) subdifferential of $q$ at $x$. Clearly, we always have ^$\widehat{\partial}q(x)\subset\partial q(x)$ by construction of these sets. Whenever $q$ is convex, both subdifferentials coincide, i.e.,
^$\widehat{\partial}q(x)=\partial q(x)=\{\eta\in\mathbb{X}\,|\,\forall y\in\operatorname{dom}q\colon\,q(y)\geq q(x)+\langle\eta,y-x\rangle\}$
is valid in this situation. By definition of the regular subdifferential, it is clear that whenever $x^{*}\in\operatorname{dom}q$ is a local minimizer of $q$, then $0\in\widehat{\partial}q(x^{*})$ hold by Fermat’s rule [22, Proposition 1.30(i)]. Thus, the inclusion $0\in\partial q(x^{*})$ is a necessary optimality condition for $x^{*}$ being a local minimizer of $q$ as well.
Since $f$ is continuously differentiabble, then for any fixed $x\in\operatorname{dom}\phi$, the sum rule
$\partial(f+\phi)(x)=\nabla f(x)+\partial\phi(x)$ (2.1)
holds, see [22, Proposition 1.30(ii)]. Fermat’s rule shows that the optimality condition
$0\in\nabla f(x^{*})+\partial\phi(x^{*})$
holds at any local minimizer $x^{*}\in\operatorname{dom}\phi$ of the composite optimization problem (Q). Any point $x^{*}\in\operatorname{dom}\phi$ satisfying this necessary optimality condition will be called an M-stationary point of (Q).
We next introduce the famous Kurdyka–Łojasiewicz property which plays a central role in this manuscript. The version of this property stated below is a generalization of the classical Kurdyka–Łojasiewicz inequality for nonsmooth functions as introduced in [2, 5, 6] and afterwards used in the local convergence analysis of several nonsmooth optimization methods, see [1, 3, 7, 9, 10, 23, 24] for a couple of examples.
Definition 2.1.
Let $g\colon\mathbb{X}\to\overline{\mathbb{R}}$ be lower semicontinuous. We say that $g$ has the KL property, where KL abbreviates Kurdyka–Łojasiewicz, at $x^{*}\in\{x\in\mathbb{X}\,|\,\partial g(x)\neq\emptyset\}$ if there exist a constant $\eta>0$, a neighborhood $U\subset\mathbb{X}$ of $x^{*}$, and a continuous concave function $\chi\colon[0,\eta]\to[0,\infty)$ which is continuously differentiable on $(0,\eta)$ and satisfies $\chi(0)=0$ as well as $\chi^{\prime}(t)>0$ for all $t\in(0,\eta)$ such that the so-called KL inequality
$\chi^{\prime}\big{(}g(x)-g(x^{*})\big{)}\operatorname{dist}\big{(}0,\partial g(x)\big{)}\geq 1$
holds for all $x\in U\cap\big{\{}x\in\mathbb{X}\,|\,g(x^{*})<g(x)<g(x^{*})+\eta\big{\}}$. The function $\chi$ from above is referred to as the desingularization function.
A popular example of the desingularization function is given by $\chi(t):=ct^{\theta}$ for $\theta\in(0,1]$ and some constant $c>0$, where the parameter $\theta$ is called the KL exponent, see [6, 19].
3 NPG with Average Line Search and Convergence Results
We first propose the nonmonotone proximal gradient method with average line search in Algorithm 3.1.
Algorithm 3.1 Nonmonotone Proximal Gradient Method with Average Line Search
0:  $\tau>1$, $0<\gamma_{\min}\leq\gamma_{\max}<\infty$, $\delta\in(0,1)$, $\frac{4}{5}\leq p_{\min}\leq 1$, $x^{0}\in\operatorname{dom}q$. Choose a positive sequence $\{p_{k}\}$ such that $p_{k}\in[p_{\min},1]$.
1:  Set $k:=0$ and $\Phi_{0}\leftarrow q(x^{0})$.
2:  while A suitable termination criterion is violated at iteration $k$ do
3:     Choose $\gamma_{k}^{0}\in[\gamma_{\min},\gamma_{\max}]$.
4:      For $i=0,1,2,\ldots$, compute a solution $x^{k,i}$ of
$\min_{x\in\mathbb{X}}\ f(x^{k})+\langle\nabla f(x^{k}),x-x^{k}\rangle+\frac{\gamma_{k,i}}{2}\|x-x^{k}\|^{2}+g(x)$ (3.1)
with $\gamma_{k,i}:=\tau^{i}\gamma_{k}^{0}$, until the acceptance criterion
$q(x^{k,i})\leq\Phi_{k}-\delta\frac{\gamma_{k,i}}{2}\|x^{k,i}-x^{k}\|^{2}$ (3.2)
holds.
5:      Denote $i_{k}:=i$ as the terminal value, and set $\gamma_{k}:=\gamma_{k,i_{k}}$ and $x^{k+1}:=x^{k,i_{k}}$.
6:      Set $\Phi_{k+1}:=(1-p_{k})\Phi_{k}+p_{k}q(x^{k+1})$.
7:     Set $k\leftarrow k+1$.
8:  end while
9:  return  $x^{k}$
In order to guarantee the convergence of Algorithm 3.1, we need the following requirement.
Assumption 3.2.
  1. (a)
    The function $q$ is bounded from below on $\operatorname{dom}g$.
  2. (b)
    The function $g$ is bounded from below by an affine function.
  3. (c)
    The function $\nabla f\colon\mathbb{X}\to\mathbb{X}$ is locally Lipschitz continuous.
Note that Assumption 3.2 (a) and (b) are the same as [18, Assumption 3.1]. Assumption 3.2 (a) guarantees that the optimization problems (Q) are solvable. Assumption 3.2 (b) implies that $g$ is coercive, which consequently implies that subproblems (3.1) always have solutions, and hence the algorithm is well-defined.
Combination of [18, Lemma 3.1] and [12, Lemma 4.1] implies that Algorithm 3.1 is well-defined. The following results are fundamental for the convergence analysis.
Proposition 3.3.
Let Assumption 3.2 hold and $\{x^{k}\}_{k\in\mathbb{N}}$ be any sequence generated by Algorithm 3.1, then
  1. (a)
    The sequence $\{\Phi_{k}\}_{k\in\mathbb{N}}$ is monotonically decreasing and
    $q(x^{k+1})+\delta\frac{(1-p_{k})\gamma_{k}}{2}\|x^{k+1}-x^{k}\|^{2}\leq\Phi_{k+1}\leq\Phi_{k}-\delta\frac{p_{k}\gamma_{k}}{2}\|x^{k+1}-x^{k}\|^{2}.$ (3.3)
  2. (b)
    Every iterate $x^{k}$ remains in the sublevel set $\mathcal{L}_{q}(x^{0})\subset\operatorname{dom}g$.
  3. (c)
    Both $\{q(x^{k})\}_{k\in\mathbb{N}}$ and $\{\Phi_{k}\}_{k\in\mathbb{N}}$ converge to some finite value $q_{*}\geq\inf q$.
  4. (d)
    $\lim_{k\to\infty}\|x^{k+1}-x^{k}\|=0$.
  5. (e)
    $\Phi_{k}\geq q(x^{k})$ holds for all $k\in\mathbb{N}$.
Proof.
Note that (a)$\sim$(d) have been observed in [12, Lemma 4.2] and [12, Lemma 4.3], so we omit the proof. It remains to consider (e).
When $k=0$, one has $\Phi_{0}=q(x^{0})$ by the initialization in Algorithm 3.1, then for any $k>0$, from (3.2), Steps 5 and 6, one has
$\Phi_{k}$ $:=(1-p_{k-1})\Phi_{k-1}+p_{k-1}q(x^{k})$
$\geq(1-p_{k-1})\left(q(x^{k})+\delta\frac{\gamma_{k-1}}{2}\|x^{k}-x^{k-1}\|^{2}\right)+p_{k-1}q(x^{k})$
$=q(x^{k})+\delta\frac{(1-p_{k-1})\gamma_{k-1}}{2}\|x^{k}-x^{k-1}\|^{2}\geq q(x^{k}).$
Hence, one has
$\Phi_{k}\geq q(x^{k})\quad\forall k\in\mathbb{N}.$
With the aid of Proposition 3.3, we will obtain the following result.
Lemma 3.4.
Let Assumption 3.2 hold, $\{x^{k}\}_{k\in\mathbb{N}}$ be any sequence generated by Algorithm 3.1, and $\bar{x}$ be an accumulation point of $\{x^{k}\}_{k\in\mathbb{N}}$. Suppose that $\{x^{k}\}_{k\in\mathcal{K}}$ is a subsequence converging to some point $\bar{x}$, then $\gamma_{k}\|x^{k+1}-x^{k}\|\to_{\mathcal{K}}0$ holds.
Proof.
If the sequence $\{\gamma_{k}\}_{\mathcal{K}}$ is bounded, the statement follows directly from Proposition 3.3 (d). It remains to consider the case where this subsequence is unbounded. Without loss of generality, we assume that $\gamma_{k}\to_{\mathcal{K}}\infty$ and then, for the trial stepsize $\hat{\gamma}_{k}:=\gamma_{k}/\tau=\tau^{i_{k}-1}\gamma_{k}^{0}$, one also has $\hat{\gamma}_{k}\to_{\mathcal{K}}\infty$, whereas the corresponding trial vector $\hat{x}^{k}:=x^{k,i_{k}-1}$ does not satisfy the acceptance criterion from (3.2), together with Proposition 3.3 (e), one has
$q(\hat{x}^{k})>\Phi_{k}-\delta\frac{\hat{\gamma}_{k}}{2}\|\hat{x}^{k}-x^{k}\|^{2}\geq q(x^{k})-\delta\frac{\hat{\gamma}_{k}}{2}\|\hat{x}^{k}-x^{k}\|^{2}\quad\forall k\in\mathcal{K}.$ (3.4)
On the other hand, since $\hat{x}^{k}$ solves the corresponding subproblem (3.1) with $\hat{\gamma}_{k}$, one has
$\langle\nabla f(x^{k}),\hat{x}^{k}-x^{k}\rangle+\frac{\hat{\gamma}_{k}}{2}\|\hat{x}^{k}-x^{k}\|^{2}+g(\hat{x}^{k})-g(x^{k})\leq 0$ (3.5)
holds for all $k\in\mathcal{K}$. We immediately obtain that $\|\hat{x}^{k}-x^{k}\|\to_{\mathcal{K}}0$ (otherwise, the left-hand side in (3.5) goes to infinity) and consequently $\hat{x}^{k}\to_{\mathcal{K}}\bar{x}$ from Assumption 3.2 (b). By Assumption 3.2 (a), using the mean-value theorem yields the existence of a point $\xi^{k}$ on the line segment connecting $x^{k}$ with $\hat{x}^{k}$ such that
$q(\hat{x}^{k})-q(x^{k})$ $=f(\hat{x}^{k})+g(\hat{x}^{k})-f(x^{k})-g(x^{k})$
$=\langle\nabla f(\xi^{k}),\hat{x}^{k}-x^{k}\rangle+g(\hat{x}^{k})-g(x^{k}).$
Substituting the resulting expression for $g(\hat{x}^{k})-g(x^{k})$ into (3.5) yields
$\langle\nabla f(x^{k})-\nabla f(\xi^{k}),\hat{x}^{k}-x^{k}\rangle+\frac{\hat{\gamma}_{k}}{2}\|\hat{x}^{k}-x^{k}\|^{2}+q(\hat{x}^{k})-q(x^{k})\leq 0$ (3.6)
for all $k\in\mathcal{K}$. Exploiting (3.4), one therefore obtains
$\frac{\hat{\gamma}_{k}}{2}\|\hat{x}^{k}-x^{k}\|^{2}$ $\leq-\langle\nabla f(x^{k})-\nabla f(\xi^{k}),\hat{x}^{k}-x^{k}\rangle+q(x^{k})-q(\hat{x}^{k})$
$<\|\nabla f(x^{k})-\nabla f(\xi^{k})\|\|\hat{x}^{k}-x^{k}\|+\delta\frac{\hat{\gamma}_{k}}{2}\|\hat{x}^{k}-x^{k}\|^{2},$
which can be rewritten as
$(1-\delta)\frac{\hat{\gamma}_{k}}{2}\|\hat{x}^{k}-x^{k}\|<\|\nabla f(x^{k})-\nabla f(\xi^{k})\|.$
Since $\xi^{k}$ is an element from the line connecting $x^{k}$ and $\hat{x}^{k}$, it follows that $\xi^{k}\to_{\mathcal{K}}\bar{x}$. Recall that $f$ is continuously differentiable, one has ${\hat{\gamma}_{k}}\|\hat{x}^{k}-x^{k}\|\to_{\mathcal{K}}0$.
Exploiting the fact that $x^{k+1}$ and $\hat{x}^{k}$ are solutions of the subproblem (3.1) with stepsize $\gamma_{k}$ and $\hat{\gamma}_{k}$, respectively, we obtain that
$\langle\nabla f(x^{k}),x^{k+1}-x^{k}\rangle$ $+\frac{\gamma_{k}}{2}\|x^{k+1}-x^{k}\|^{2}+g(x^{k+1})$
$\leq\langle\nabla f(x^{k}),\hat{x}^{k}-x^{k}\rangle+\frac{\gamma_{k}}{2}\|\hat{x}^{k}-x^{k}\|^{2}+g(\hat{x}^{k})$
and
$\langle\nabla f(x^{k}),\hat{x}^{k}-x^{k}\rangle$ $+\frac{\hat{\gamma}_{k}}{2}\|\hat{x}^{k}-x^{k}\|^{2}+g(\hat{x}^{k})$
$\leq\langle\nabla f(x^{k}),x^{k+1}-x^{k}\rangle+\frac{\hat{\gamma}_{k}}{2}\|x^{k+1}-x^{k}\|^{2}+g(x^{k+1}).$
Adding the two inequalities together, then the fact that $\gamma_{k}=\tau\hat{\gamma}_{k}$ implies $\|x^{k+1}-x^{k}\|\leq\|\hat{x}^{k}-x^{k}\|$, and therefore,
$\gamma_{k}\|x^{k+1}-x^{k}\|=\tau\hat{\gamma}_{k}\|x^{k+1}-x^{k}\|\leq\tau\hat{\gamma}_{k}\|\hat{x}^{k}-x^{k}\|\to_{\mathcal{K}}0.$
This completes the proof. ∎
By means of the proof of Lemma 3.4 and the technique of [15, Lemma 4.1], we immediately have the following desired result, i.e., around the accumulation point $\bar{x}$ of $\{x^{k}\}_{k\in\mathbb{N}}$, the associated stepsize sequence remains (uniformly) bounded. This plays a central role in the convergence of entire sequence. We omit the proof because it is highly similar with [17, Lemma 4.1].
Lemma 3.5.
Let Assumption 3.2 hold, $\{x^{k}\}_{k\in\mathbb{N}}$ be any sequence generated by Algorithm 3.1, and $\bar{x}$ is an accumulation point of $\{x^{k}\}_{k\in\mathbb{N}}$. Then, for any $\rho>0$, there is a constant $\bar{\gamma}_{\rho}>0$ (usually depending on $\rho$) such that $\gamma_{k}\leq\bar{\gamma}_{\rho}$ for all $k\in\mathbb{N}$ such that $x^{k}\in B_{\rho}(\bar{x})$.
The following result provides the main global (subsequential) convergence of Algorithm 3.1.
Theorem 3.6.
Let Assumption 3.2 hold. Then each accumulation point $\bar{x}$ of a sequence $\{x^{k}\}_{k\in\mathbb{N}}$ generated by Algorithm 3.1 is a stationary point of (Q), and $q_{*}=q(\bar{x})$.
Proof.
Let $\{x^{k}\}_{\mathcal{K}}$ be a subsequence converging to $\bar{x}$. Recall again that $x^{k+1}$ is a solution of the subproblem (3.1), hence one has
$0\in\nabla f(x^{k})+\gamma_{k}(x^{k+1}-x^{k})+\partial g(x^{k+1})\quad\forall k\in\mathbb{N}.$
If we have $g(x^{k+1})\to_{\mathcal{K}}g(\bar{x})$, taking $k\to_{\mathcal{K}}\infty$ to the above optimality condition, then we have $0\in\nabla f(\bar{x})+\partial g(\bar{x})$ from Lemma 3.4 and the continuity of $\nabla f$, which means that $\bar{x}$ is imediately a stationary point of (Q) and $q_{*}=q(\bar{x})$ holds from Proposition 3.3 (c) and the continuity of $f$. So, it remains to prove $g(x^{k+1})\to_{\mathcal{K}}g(\bar{x})$.
From the fact that $g$ is lower semicontinuous, one then has
$g(\bar{x})\leq\liminf_{k\to_{\mathcal{K}}\infty}g(x^{k+1})\leq\limsup_{k\to_{\mathcal{K}}\infty}g(x^{k+1}).$ (3.7)
Since again $x^{k+1}$ is a solution of the subproblem (3.1) with stepsize $\gamma_{k}$, hence one has
$\langle\nabla f(x^{k}),x^{k+1}-x^{k}\rangle+\frac{\gamma_{k}}{2}\|x^{k+1}-x^{k}\|^{2}+g(x^{k+1})\leq\langle\nabla f(x^{k}),\bar{x}-x^{k}\rangle+\frac{\gamma_{k}}{2}\|\bar{x}-x^{k}\|^{2}+g(\bar{x})$ (3.8)
for all $k\in\mathbb{N}$. Recall again that $\bar{x}$ is an accumulation point of $\{x^{k}\}_{k\in\mathbb{N}}$, hence $\gamma_{k}$ is finite for sufficiently large $k\in\mathcal{K}$ from Lemma 3.5. Taking $k\to_{\mathcal{K}}\infty$ to (3.8), one has $\lim\sup_{k\to_{\mathcal{K}}\infty}g(x^{k+1})\leq g(\bar{x})$ from the fact that $x^{k}\to_{\mathcal{K}}\bar{x}$, the continuity of $\nabla f$, Proposition 3.3 (d), and Lemma 3.4. In view of (3.7), one has $g(x^{k+1})\to_{\mathcal{K}}g(\bar{x})$, this completes the proof. ∎
Theorem 3.6 actually obtained the subsequential convergence, we next aim to give the convergence of the entire sequence. Let sufficiently small $\eta>0$ be the corresponding constant from the definition of the associated desingularization function $\chi$. In view of Proposition 3.3 (d), one can find a sufficiently large index $\hat{k}\in\mathbb{N}$ such that
$\sup_{k\geq\hat{k}}\|x^{k+1}-x^{k}\|\leq\eta.$ (3.9)
We denote
$\rho:=\eta+\frac{1}{2},$ (3.10)
as well as the compact set
$C_{\rho}:=B_{\rho}(\bar{x}).$ (3.11)
Finally, throughout the section, let $L_{\rho}>0$ be a (global) Lipschitz constant of $\nabla f$ on $C_{\rho}$ from (3.11). In view of Lemma 3.5, one has
$\gamma_{k}\leq\bar{\gamma}_{\rho}\quad\forall x^{k}\in C_{\rho}$ (3.12)
with some suitable upper bound $\bar{\gamma}_{\rho}>0$ (depending on our choice of $\rho$ from (3.10)). Since $p_{\min}>\frac{4}{5}$, then we denote
$l:=\frac{1}{2}-\sqrt{\frac{1-p_{\min}}{p_{\min}}}>0.$ (3.13)
In view of Proposition 3.3 (e), we know that $\Phi_{k}\geq q(x^{k})$ for all $k\in\mathbb{N}$, however the quantitative relationship between $\Phi_{k+1}$ and $q(x^{k})$ is uncertain, so in the following, we define
$S:=\{k\in\mathbb{N}\,|\,q(x^{k})-\Phi_{k+1}\leq\frac{\mu}{2}\|x^{k+1}-x^{k}\|^{2}\},$ (3.14)
where $\mu\in(0,\frac{1}{2}\delta{p_{\min}\gamma_{\min}}]$ is a constant. Also define $\overline{S}:=\mathbb{N}\setminus S$.
Lemma 3.7.
Let Assumption 3.2 hold and $\{x^{k}\}_{k\in\mathbb{N}}$ be any sequence generated by Algorithm 3.1, and $\bar{x}$ be an accumulation point of $\{x^{k}\}_{k\in\mathbb{N}}$. Suppose that $\{x^{k}\}_{k\in\mathcal{K}}$ is a subsequence converging to $\bar{x}$, and that $q$ has the KL property at $\bar{x}$ with desingularization function $\chi$. Then there is a sufficiently large constant $k_{0}\in\mathcal{K}\cap\overline{S}$ such that the corresponding constant
$\alpha:=\|x^{k_{0}}-\bar{x}\|+\frac{1}{l}\sqrt{\frac{2}{\delta p_{\min}\gamma_{\min}}}\sqrt{\Phi_{k_{0}}-q(\bar{x})}+\frac{\bar{\gamma}_{\rho}+L_{\rho}}{2l}\chi\big{(}\Phi_{k_{0}}-q(\bar{x})\big{)}.$ (3.15)
satisfies $\alpha<\frac{1}{2}$, where $\rho>0$, $\bar{\gamma}_{\rho}>0$, and $l>0$ are the constants defined in (3.10), (3.12), and (3.13), respectively, while $L_{\rho}>0$ is a Lipschitz constant of $\nabla f$ on $C_{\rho}$ from (3.11) and $\delta>0$ as well as $\gamma_{\min}>0$ are the parameters from Algorithm 3.1.
Proof.
From $x^{k}\to_{\mathcal{K}}\bar{x}$, then there exists a sufficiently large $k_{0}\in\mathcal{K}$ satisfying $\Phi_{k_{0}}-q(\bar{x})\geq 0$ sufficiently small deduced by Proposition 3.3 and Theorem 3.6. Recall that that $q$ satisfies KL property at $\bar{x}$, the corresponding disingularizatioon function is denoted as $\chi$. For such $\chi$, one has $\chi(0)=0$ and $\chi$ is continuous and increasing on its domain, therefore one knows $\chi\big{(}\Phi_{k_{0}}-q(\bar{x})\big{)}\geq 0$ sufficiently small. Hence, each summand on the right-hand side (3.15) by taking an index $k_{0}\in\mathcal{K}$ sufficiently large yields $\alpha<\frac{1}{2}$. ∎
Using these notations, we get the following result. We omit the proof, since it is highly similar with [17, Lemma 4.3].
Lemma 3.8.
Let Assumption 3.2 hold, $\{x^{k}\}_{k\in\mathbb{N}}$ be any sequence generated by Algorithm 3.1, and $\bar{x}$ be an accumulation point of $\{x^{k}\}_{k\in\mathbb{N}}$. Suppose that $\{x^{k}\}_{k\in\mathcal{K}}$ is a subsequence converging to $\bar{x}$, and that $q$ has the KL property at $\bar{x}$ with desingularization function $\chi$. Then
$\operatorname{dist}\big{(}0,\partial q(x^{k+1})\big{)}\leq\big{(}\bar{\gamma}_{\rho}+L_{\rho}\big{)}\|x^{k+1}-x^{k}\|$
holds for all sufficiently large $k\geq\hat{k}$ such that $x^{k}\in B_{\alpha}(\bar{x})$, where $\alpha$ denotes the constant from (3.15), $\bar{\gamma}_{\rho}>0$ is the constant from (3.12), and $L_{\rho}>0$ is the Lipschitz constant of $\nabla f$ on $C_{\rho}$ from (3.11).
By employing the above results, the following theorem demonstrates that the whole sequence $\{x^{k}\}_{k\in\mathbb{N}}$ generated by Algorithm 3.1 is convergent to its accumulation point $\bar{x}$, provided that $q$ satisfies the KL property at $\bar{x}$.
Theorem 3.9.
Let Assumption 3.2 hold, $\{x^{k}\}_{k\in\mathbb{N}}$ be any sequence generated by Algorithm 3.1, and $\bar{x}$ be an accumulation point of $\{x^{k}\}_{k\in\mathbb{N}}$. Suppose that $\{x^{k}\}_{k\in\mathcal{K}}$ is a subsequence converging to $\bar{x}$, and that $q$ has the KL property at $\bar{x}$. Then the entire sequence $\{x^{k}\}_{k\in\mathbb{N}}$ converges to $\bar{x}$.
Proof.
By Proposition 3.3 (a) and (c), one knows that the whole sequence $\{\Phi_{k}\}_{k\in\mathbb{N}}$ is monotonically decreasing and converges to $q(\bar{x})$ from Theorem 3.6. This implies that $\Phi_{k}\geq q(\bar{x})$ holds for all $k\in\mathbb{N}$. If $\Phi_{k}=q(\bar{x})$ holds for some index $k\in\mathbb{N}$, which, by monotonicity, implies that $\Phi_{k+1}=q(\bar{x})$. Let us now recall Step 6 in Algorithm 3.1, one has
$q(x^{k+1})=\frac{1}{p_{k}}(\Phi_{k+1}-\Phi_{k})+\Phi_{k}\quad\forall k\in\mathbb{N}.$ (3.16)
Hence, the acceptance criterion (3.2) implies that
$0\leq\|x^{k+1}-x^{k}\|\leq\sqrt{\frac{2\big{(}\Phi_{k}-q(x^{k+1})\big{)}}{\delta\gamma_{\min}}}=\sqrt{\frac{2(\Phi_{k}-\Phi_{k+1})}{p_{k}\delta\gamma_{\min}}}=0,$
which says that $x^{k+1}=x^{k}$ holds. Since, by assumption, the subsequence $\{x^{k}\}_{\mathcal{K}}$ converges to $\bar{x}$, this implies that $x^{k}=\bar{x}$ for all $k\in\mathbb{N}$ sufficiently large. In particular, one has convergence of the entire (eventually constant) sequence $\{x^{k}\}_{k\in\mathbb{N}}$ to $\bar{x}$ in this situation.
It remains to consider the case where $\Phi_{k}>q(\bar{x})$ for all $k\in\mathbb{N}$. Let $\alpha$ be the constant from (3.15), and $k_{0}\in\mathcal{K}$ be the corresponding iteration index which is used in the definition of $\alpha$, see Lemma 3.7. One then has $0<\Phi_{k}-q(\bar{x})\leq\Phi_{k_{0}}-q(\bar{x})$ for all $k\geq k_{0}$. Without loss of generality, we may also assume that $k_{0}\geq\hat{k}$ defined by (3.9) and that $k_{0}$ is sufficiently large to satisfy
$\Phi_{k_{0}}<q(\bar{x})+\eta.$ (3.17)
Let $\chi\colon[0,\eta]\to[0,\infty)$ be the desingularization function which comes along with the validity of the KL property of $q$ at $\bar{x}$. Due to $\chi(0)=0$ and $\chi^{\prime}(t)>0$ for all $t\in(0,\eta)$, one obtains
$\chi\big{(}\Phi_{k}-q(\bar{x})\big{)}\geq 0\quad\forall k\geq k_{0}.$ (3.18)
We now claim that the following two statements hold for all $k\geq k_{0}$:
  1. (a)
    $x^{k}\in B_{\alpha}(\bar{x})$,
  2. (b)
    $\|x^{k_{0}}-\bar{x}\|+\sum_{i=k_{0}}^{k}\|x^{i+1}-x^{i}\|\leq\alpha$, which is equivalent to
    $\sum_{i=k_{0}}^{k}\|x^{i+1}-x^{i}\|$ $\leq\frac{1}{l}\sqrt{\frac{2}{\delta p_{\min}\gamma_{\min}}}\sqrt{\Phi_{k_{0}}-q(\bar{x})}+\frac{\bar{\gamma}_{\rho}+L_{\rho}}{2l}\chi\big{(}\Phi_{k_{0}}-q(\bar{x})\big{)}.$ (3.19)
We verify these two statements jointly by induction. For $k=k_{0}$, statement (a) holds simply by the definition of $\alpha$ in (3.15). Meanwhile, (3.2) and (3.16) implies
$\|x^{k+1}-x^{k}\|$ $\leq\sqrt{\frac{2\big{(}\Phi_{k}-q(x^{k+1})\big{)}}{\delta\gamma_{\min}}}=\sqrt{\frac{2(\Phi_{k}-\Phi_{k+1})}{p_{k}\delta\gamma_{\min}}}\leq\sqrt{\frac{2\big{(}\Phi_{k}-\Phi_{k+1}\big{)}}{\delta p_{\min}\gamma_{\min}}}$ (3.20)
holds for all $k\in\mathbb{N}$. Recall the fact that $\{\Phi_{k}\}_{k\in\mathbb{N}}$ is decreasing and bounded below by $q(\bar{x})$, then (3.20) implies that
$\|x^{k+1}-x^{k}\|\leq\sqrt{\frac{2\big{(}\Phi_{k_{0}}-q(\bar{x})\big{)}}{\delta p_{\min}\gamma_{\min}}}\quad\forall k\geq k_{0},$
hence (3.19) holds for $k=k_{0}$ due to $11/l>$ 1. Suppose that both statements hold for some $k\geq k_{0}$. Using the triangle inequality, the induction hypothesis, and the definition of $\alpha$, one obtains
$\|x^{k+1}-\bar{x}\|\leq\sum_{i=k_{0}}^{k}\|x^{i+1}-x^{i}\|+\|x^{k_{0}}-\bar{x}\|\leq\alpha,$
i.e., statement (a) holds for $k+1$ in place of $k$. The verification of the induction step for (b) is more involved.
By employing the index set $S$ defined in (3.14), we consider the following two cases for sufficiently large $k\geq k_{0}$.
Case 1: $k\in S$. We know that
$q(x^{k})-\Phi_{k+1}\leq\frac{\mu}{2}\|x^{k+1}-x^{k}\|^{2}\quad\forall k\in S.$
Then (3.3), (3.14), and (3.16) imply
$\Phi_{k}-\Phi_{k+1}$ $\geq\delta\frac{p_{\min}\gamma_{\min}}{2}\|x^{k+1}-x^{k}\|^{2}\geq\frac{\delta p_{\min}\gamma_{\min}}{\mu}\big{(}q(x^{k})-\Phi_{k+1}\big{)}\geq 2\big{(}q(x^{k})-\Phi_{k+1}\big{)}$
$=2\left(\frac{1}{p_{k-1}}\big{(}\Phi_{k}-\Phi_{k-1}\big{)}+\Phi_{k-1}-\Phi_{k+1}\right)$
$=2\left(\frac{1}{p_{k-1}}-1\right)\big{(}\Phi_{k}-\Phi_{k-1}\big{)}+2(\Phi_{k}-\Phi_{k+1})\quad\forall S\ni k\geq k_{0}+1,$
which yields
$\Phi_{k}-\Phi_{k+1}\leq 2\left(\frac{1-p_{k-1}}{p_{k-1}}\right)\big{(}\Phi_{k-1}-\Phi_{k}\big{)}\leq 2\left(\frac{1-p_{\min}}{p_{\min}}\right)\big{(}\Phi_{k-1}-\Phi_{k}\big{)}\quad\forall S\ni k\geq k_{0}+1.$
Due to $k_{0}\in\overline{S}$, summation yields that
$\sum_{{S\ni i=k_{0}}}^{k+1}\sqrt{\Phi_{i}-\Phi_{i+1}}$ $=\sum_{{S\ni i=k_{0}+1}}^{k+1}\sqrt{\Phi_{i}-\Phi_{i+1}}$ (3.21)
$\leq\sqrt{\frac{2(1-p_{\min})}{p_{\min}}}\sum_{{S\ni i=k_{0}+1}}^{k+1}\sqrt{\Phi_{i-1}-\Phi_{i}}.$
Case 2: $k\in\overline{S}$. For all $i$ satisfying $i\geq k_{0}$ and $i\in\overline{S}$, then one has $q(x^{i})>\Phi_{i+1}$, and hence $q(x^{i})\geq q(x^{i+1})$ holds. From Proposition 3.3 (e) and (3.17), one has that
$q(\bar{x})<\Phi_{i+1}<q(x^{i})\leq\Phi_{i}\leq\Phi_{k_{0}}<q(\bar{x})+\eta\quad\forall i\geq k_{0}\ \text{and}\ i\in\overline{S}.$ (3.22)
Since $q$ has the KL property at $\bar{x}$, one has
$\chi^{\prime}\big{(}q(x^{i})-q(\bar{x})\big{)}\operatorname{dist}\big{(}0,\partial q(x^{i})\big{)}\geq 1\quad\forall i\geq k_{0}\ \text{and}\ i\in\overline{S}.$ (3.23)
Since $x^{i}\in B_{\alpha}(\bar{x})$ holds for all $i\in\{k_{0},k_{0}+1,\ldots,k\}$ by our induction hypothesis, we can apply Lemma 3.8 and obtain
$\operatorname{dist}\big{(}0,\partial q(x^{i+1})\big{)}\leq(\bar{\gamma}_{\rho}+L_{\rho})\|x^{i+1}-x^{i}\|\quad\forall i\in k_{0},\ldots,k,$
hence, one has
$\operatorname{dist}\big{(}0,\partial q(x^{i})\big{)}\leq(\bar{\gamma}_{\rho}+L_{\rho})\|x^{i}-x^{i-1}\|\quad\forall i\in\{k_{0}+1,\ldots,k+1\}\cap\overline{S}.$ (3.24)
Recall that $\chi$ is increasing and concave on $[0,\eta)$, $\chi^{\prime}(t)>0$ for all $t>0$, and $\chi(0)=0$, then one has
$\chi\big{(}\Phi_{k}-q(\bar{x})\big{)}-\chi\big{(}\Phi_{k+1}-q(\bar{x})\big{)}$ $\geq\chi\big{(}q(x^{k})-q(\bar{x})\big{)}-\chi\big{(}\Phi_{k+1}-q(\bar{x})\big{)}$ (3.25)
$\geq\chi^{\prime}\big{(}q(x^{k})-q(\bar{x})\big{)}\big{(}q(x^{k})-\Phi_{k+1}\big{)}$
$\geq\frac{q(x^{k})-\Phi_{k+1}}{(\bar{\gamma}_{\rho}+L_{\rho})\|x^{k}-x^{k-1}\|}$
$\geq\sqrt{\frac{2}{\delta p_{\min}\gamma_{\min}}}\frac{q(x^{k})-\Phi_{k+1}}{(\bar{\gamma}_{\rho}+L_{\rho})\sqrt{\Phi_{k-1}-\Phi_{k}}}$
$:=c\frac{q(x^{k})-\Phi_{k+1}}{\sqrt{\Phi_{k-1}-\Phi_{k}}}$
for all $\{k_{0}+1,\ldots,k+1\}\cap\overline{S}$, where $c:=\sqrt{\frac{2}{\delta p_{\min}\gamma_{\min}}}\frac{1}{(\bar{\gamma}_{\rho}+L_{\rho})}$. Meanwhile, (3.2) means that $q(x^{k+1})\leq\Phi_{k}$, it follows that
$\Phi_{k+1}:=(1-p_{k})\Phi_{k}+p_{k}q(x^{k+1})\leq(1-p_{\min})\Phi_{k}+p_{\min}q(x^{k+1})\quad\forall k\in\mathbb{N}.$ (3.26)
Denote $\Delta_{i,j}=\chi\big{(}\Phi_{i}-q(\bar{x})\big{)}-\chi\big{(}\Phi_{j}-q(\bar{x})\big{)}$ for convenience, (3.25) and (3.26) yield that
$\Phi_{k}-\Phi_{k+1}$ $\leq(1-p_{\min})\Phi_{k-1}+p_{\min}q(x^{k})-\Phi_{k+1}$
$\leq(1-p_{\min})\big{(}\Phi_{k-1}-\Phi_{k+1}\big{)}+p_{\min}\big{(}q(x^{k})-\Phi_{k+1}\big{)}$
$\leq(1-p_{\min})\big{(}\Phi_{k-1}-\Phi_{k+1}\big{)}+\frac{p_{\min}}{c}\Delta_{k,k+1}\sqrt{\Phi_{k-1}-\Phi_{k}}$
$\leq(1-p_{\min})\left(\big{(}\Phi_{k-1}-\Phi_{k}\big{)}+\big{(}\Phi_{k}-\Phi_{k+1}\big{)}\right)+\frac{p_{\min}}{c}\Delta_{k,k+1}\sqrt{\Phi_{k-1}-\Phi_{k}},$
which implies that
$p_{\min}\big{(}\Phi_{k}-\Phi_{k+1}\big{)}\leq(1-p_{\min})\big{(}\Phi_{k-1}-\Phi_{k}\big{)}+\frac{p_{\min}}{c}\Delta_{k,k+1}\sqrt{\Phi_{k-1}-\Phi_{k}}.$
Using $\sqrt{a+b}\leq\sqrt{a}+\sqrt{b}$ and $2\sqrt{ab}\leq a+b$ for all $a,b\geq 0$, we obtain
$\sqrt{p_{\min}}\sqrt{\Phi_{k}-\Phi_{k+1}}$ $\leq\sqrt{1-p_{\min}}\sqrt{\Phi_{k-1}-\Phi_{k}}+\sqrt{p_{\min}}\sqrt{\frac{1}{c}\Delta_{k,k+1}\sqrt{\Phi_{k-1}-\Phi_{k}}}$
$\leq\sqrt{1-p_{\min}}\sqrt{\Phi_{k-1}-\Phi_{k}}+\frac{\sqrt{p_{\min}}}{2}\left(\frac{1}{c}\Delta_{k,k+1}+\sqrt{\Phi_{k-1}-\Phi_{k}}\right),$
which yields that
$\sqrt{p_{\min}}\sqrt{\Phi_{k}-\Phi_{k+1}}\leq\left(\frac{\sqrt{p_{\min}}}{2}+\sqrt{1-p_{\min}}\right)\sqrt{\Phi_{k-1}-\Phi_{k}}+\frac{\sqrt{p_{\min}}}{2c}\Delta_{k,k+1}$
for all $\{k_{0}+1,\ldots,k+1\}\cap\overline{S}$. Summation yields that
$\sum_{\overline{S}\ni i=k_{0}+1}^{k+1}\sqrt{\Phi_{i}-\Phi_{i+1}}$ $\leq\sum_{\overline{S}\ni i=k_{0}+1}^{k+1}\left(\frac{1}{2}+\sqrt{\frac{1-p_{\min}}{p_{\min}}}\right)\sqrt{\Phi_{i-1}-\Phi_{i}}+\frac{1}{2c}\Delta_{i,i+1}$
$\leq\sum_{\overline{S}\ni i=k_{0}+1}^{k+1}\left(\frac{1}{2}+\sqrt{\frac{1-p_{\min}}{p_{\min}}}\right)\sqrt{\Phi_{i-1}-\Phi_{i}}+\sum_{i=k_{0}+1}^{k+1}\frac{1}{2c}\Delta_{i,i+1}$
$\leq\sum_{\overline{S}\ni i=k_{0}+1}^{k+1}\left(\frac{1}{2}+\sqrt{\frac{1-p_{\min}}{p_{\min}}}\right)\sqrt{\Phi_{i-1}-\Phi_{i}}+\frac{1}{2c}\chi\big{(}\Phi_{k_{0}}-q(\bar{x})\big{)}.$
Due to $k_{0}\in\overline{S}$, we then obtain
$\sum_{\overline{S}\ni i=k_{0}}^{k+1}\sqrt{\Phi_{i}-\Phi_{i+1}}$ $\leq\sqrt{\Phi_{k_{0}}-\Phi_{k_{0}+1}}+\sum_{\overline{S}\ni i=k_{0}+1}^{k+1}\sqrt{\Phi_{i}-\Phi_{i+1}}$ (3.27)
$\leq\sqrt{\Phi_{k_{0}}-q(\bar{x})}+\frac{1}{2c}\chi\big{(}\Phi_{k_{0}}-q(\bar{x})\big{)}$
$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ +\sum_{\overline{S}\ni i=k_{0}+1}^{k+1}\left(\frac{1}{2}+\sqrt{\frac{1-p_{\min}}{p_{\min}}}\right)\sqrt{\Phi_{i-1}-\Phi_{i}}.$
Since $p_{\min}>\frac{4}{5}$, then $\sqrt{\frac{2(1-p_{\min})}{p_{\min}}}<\frac{1}{2}+\sqrt{\frac{1-p_{\min}}{p_{\min}}}$ holds. Adding (3.21) and (3.27) yields that
$\sum_{i=k_{0}}^{k+1}\sqrt{\Phi_{i}-\Phi_{i+1}}$ $\leq\sqrt{\Phi_{k_{0}}-q(\bar{x})}+\frac{1}{2c}\chi\big{(}\Phi_{k_{0}}-q(\bar{x})\big{)}$
$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ +\left(\frac{1}{2}+\sqrt{\frac{1-p_{\min}}{p_{\min}}}\right)\sum_{i=k_{0}+1}^{k+1}\sqrt{\Phi_{i-1}-\Phi_{i}}.$
Recall again (3.13), we obtain
$\sum_{i=k_{0}}^{k+1}\sqrt{\Phi_{i}-\Phi_{i+1}}\leq\frac{1}{l}\sqrt{\Phi_{k_{0}}-q(\bar{x})}+\frac{1}{2cl}\chi\big{(}\Phi_{k_{0}}-q(\bar{x})\big{)}.$
Then, together with (3.20), we have
$\sum_{i=k_{0}}^{k+1}\|x^{i+1}-x^{i}\|$ $\leq\sum_{i=k_{0}}^{k+1}\sqrt{\frac{2\big{(}\Phi_{i}-\Phi_{i+1}\big{)}}{\delta p_{\min}\gamma_{\min}}}$
$\leq\frac{1}{l}\sqrt{\frac{2}{\delta p_{\min}\gamma_{\min}}}\sqrt{\Phi_{k_{0}}-q(\bar{x})}+\frac{\bar{\gamma}_{\rho}+L_{\rho}}{2l}\chi\big{(}\Phi_{k_{0}}-q(\bar{x})\big{)}.$
Hence, statement (b) holds for $k+1$ in place of $k$, and this completes the induction.
In particular, it follows from (a) that $x^{k}\in B_{\alpha}(\bar{x})$ for all $k\geq k_{0}$. And taking $k\to\infty$ in (3.19) therefore shows that $\{x^{k}\}_{k\in\mathbb{N}}$ is a Cauchy sequence and, thus, convergent. Since one already knows that $\bar{x}$ is an accumulation point, it follows that the entire sequence $\{x^{k}\}_{k\in\mathbb{N}}$ converges to $\bar{x}$. ∎
Note that if all index $k\geq k_{0}$ belong to $S$ or if $p_{k}:=1$, then Algorithm 3.1 degenerates into monotone proximal gradient, see [18, Algorithm 3.1] or [15, Algorithm 3.1], and the corresponding rate-of-convergence result was obtained in [14]. So, we in the following state the rate-of-convergence result of the nonmonotone proximal gradient methods (in general case), readers may find more details for the proof in [17, Theorem 4.6].
Theorem 3.10.
Let Assumption 3.2 hold and $\{x^{k}\}_{k\in\mathbb{N}}$ be any sequence generated by Algorithm 3.1. Suppose that $\{x^{k}\}_{k\in\mathcal{K}}$ is a subsequence converging to some limit point $\bar{x}$, and that $q$ has the KL property at $\bar{x}$. Then the entire sequence $\{x^{k}\}_{k\in\mathbb{N}}$ converges to $\bar{x}$, and if the corresponding desingularization function has the form $\chi(t)=\kappa t^{\theta}$ for some $kappa>0$ and $\theta\in(0,1]$, then the following statements hold:
  1. (a)
    if $\theta=1$, then the sequences $\{\Phi_{k}\}_{k\in\mathbb{N}}$ and $\{x^{k}\}_{k\in\mathbb{N}}$ converge in a finite number of steps to $q(\bar{x})$ and $\bar{x}$, respectively.
  2. (b)
    if $\theta\in[\frac{1}{2},1)$, then $\{\Phi_{k}\}$ Q-linearly convergent to $q(\bar{x})$, and $\{x^{k}\}$ R-linearly convergent to $\bar{x}$.
  3. (c)
    if $\theta\in(0,\frac{1}{2})$, then there exist some positive constants $\eta_{1}$ and $\eta_{2}$ such that
    $\Phi_{k}-q(\bar{x})$ $\leq\eta_{1}k^{-\frac{1}{1-2\theta}},$
    $\|x^{k}-\bar{x}\|$ $\leq\eta_{2}{k}^{-\frac{\theta}{1-2\theta}}$
    for sufficiently large $k\geq k_{0}$.
4 NPG with Max Line Search and Convergence Analysis
This section focuses on another type of nonmonotone proximal gradient method, where the line search is chosen as max-type rule. Let us now introduce the algorithm which is exactly from [18, Algorithm 4.1].
Algorithm 4.1 Nonmonotone Proximal Gradient Method with Max Line Search
0:  $\tau>1$, $0<\gamma_{\min}\leq\gamma_{\max}<\infty$, $m\in\mathbb{N}$, $\delta\in(0,1)$, $x^{0}\in\operatorname{dom}q$.
1:  Set $k:=0$.
2:  while A suitable termination criterion is violated at iteration $k$ do
3:     Set $m_{k}:=\min\{k,m\}$ and choose $\gamma_{k}^{0}\in[\gamma_{\min},\gamma_{\max}]$.
4:      For $i=0,1,2,\ldots$, compute a solution $x^{k,i}$ of
$\min_{x}\ f(x^{k})+\langle\nabla f(x^{k}),x-x^{k}\rangle+\frac{\gamma_{k,i}}{2}\|x-x^{k}\|^{2}+g(x),\quad x\in\mathbb{X}$ (4.1)
with $\gamma_{k,i}:=\tau^{i}\gamma_{k}^{0}$, until the acceptance criterion
$q(x^{k,i})\leq\max_{j=0,1,\ldots,m_{k}}q(x^{k-j})-\delta\frac{\gamma_{k,i}}{2}\|x^{k,i}-x^{k}\|^{2}$ (4.2)
holds.
5:      Denote by $i_{k}:=i$ the terminal value, and set $\gamma_{k}:=\gamma_{k,i_{k}}$ and $x^{k+1}:=x^{k,i_{k}}$.
6:     Set $k\leftarrow k+1$.
7:  end while
8:  return  $x^{k}$
Throughout the section, let us denote $l(k)\in\{k-m_{k},\ldots,k\}$ an index satisfying
$q(x^{l(k)})=\max_{j=0,1,\ldots,m_{k}}q(x^{k-j})$
for each $k\in\mathbb{N}$. In order to ensure the convergence of Algorithm 4.1, we need both Assumption 3.2 and
Assumption 4.2.
The function $q$ is continuous on $\operatorname{dom}q$.
Note that Assumption 3.2 (b) and Assumption 4.2 imply that $q$ is uniformly continuous on the corresponding sublevel $\mathcal{L}_{q}(x^{0})$. The requirement of uniformly continuity about objective function plays an essential role in the context of nonmonotone line search rules with maximum taste [13]. We here recall the results again for the convenience of readers.
Proposition 4.3.
Let Assumption 3.2 and Assumption 4.2 hold and let $\{x^{k}\}$ be any sequence generated by Algorithm 4.1, then
  1. (a)
    the sequence $\{q(x^{l(k)})\}$ is monotonically decreasing,
  2. (b)
    the sequences $\{x^{k}\}$, $\{x^{l(k)}\}$ $\subset\mathcal{L}_{q}(x^{0})\subset\operatorname{dom}g$,
  3. (c)
    $\lim_{k\to\infty}\|x^{k+1}-x^{k}\|=0$,
  4. (d)
    suppose $\bar{x}$ is an accumulation point of $\{x^{k}\}$ such that $x^{k}\to_{\mathcal{K}}\bar{x}$ holds. Then $\bar{x}$ is an M-stationary point of (Q), and $\gamma_{k}\|x^{k+1}-x^{k}\|\to_{\mathcal{K}}0$ is valid.
With the aid of Proposition 4.3 (d) as well as the techniques in Lemma 3.5, we can obtain the following results.
Lemma 4.4.
Let Assumption 3.2 and Assumption 4.2 hold, let $\{x^{k}\}$ be any sequence generated by Algorithm 4.1, $\bar{x}$ is an accumulation point of $\{x^{k}\}$ and consequently let $\{x^{k}\}_{\mathcal{K}}$ be a subsequence converging to some point $\bar{x}$. Then, for any $\rho>0$, there is a constant $\bar{\gamma}_{\rho}>0$ (usually depending on $\rho$) such that $\gamma_{k}\leq\bar{\gamma}_{\rho}$ for all $k\in\mathbb{N}$ such that $x^{k}\in B_{\rho}(\bar{x})$.
We now discuss the convergence properties of the sequence generated by Algorithm 4.1 in the presence of KL property of $q$ at the mentioned accumulation point $\bar{x}$. Let sufficiently small $\eta>0$ be the corresponding constant from the definition of the associated desingularization function $\chi$. In view of Lemma 4.4, we can find a sufficiently large index $\hat{k}\in\mathbb{N}$ such that
$\sup_{k\geq\hat{k}-m-1}\|x^{k+1}-x^{k}\|\leq\eta.$ (4.3)
Let us now fix a constant $\mu\in(0,\delta\gamma_{\min})$, and define
$K:=\big{\{}k\in\mathbb{N}\,|\,q(x^{l(k+1)})-q(x^{k+1})>\frac{\mu}{2}\|x^{k+1}-x^{k}\|^{2}\big{\}}$ (4.4)
and $\overline{K}:=\mathbb{N}\setminus K$. Define
$\rho:=(m+1)\eta+\frac{1}{2},$ (4.5)
as well as the compact set
$C_{\rho}:=B_{2\rho}(\bar{x})\cap\mathcal{L}_{q}(x^{0}).$ (4.6)
Consequently, let $L_{\rho}>0$ be a (global) Lipschitz constant of $\nabla f$ on $C_{\rho}$ from (4.6). In view of Lemma 4.4, let $\bar{\gamma}_{\rho}>0$ be satisfying
$\gamma_{k}\leq\bar{\gamma}_{\rho}\quad\forall x^{k}\in C_{\rho}.$ (4.7)
Thanks to the above notation, we can find a new neighborhood centered at the accumulation point, whose radium is very small which is mainly used to to ensure the convergence of the whole sequence under the assumption of KL property of $q$, we now introduce such radium.
Lemma 4.5.
Let Assumption 3.2 and Assumption 4.2 hold, and let $\{x^{k}\}$ be any sequence generated by Algorithm 4.1. Suppose that $\{x^{k}\}_{\mathcal{K}}$ is a subsequence converging to some limit point $\bar{x}$, and that $q$ has the KL property at $\bar{x}$ with desingularization function $\chi$. Then there is a sufficiently large constant $k_{0}\in\mathcal{K}\cap\overline{K}$ such that $k_{0}-m-1$ is also sufficiently large, where $K$ is defined by (4.4), then the corresponding constant is defined as
$\alpha$ $:=\|x^{k_{0}}-\bar{x}\|$ (4.8)
$\leq\left(\frac{8m(\bar{\gamma}_{\rho}+L_{\rho})}{3\delta\gamma_{\min}}\sqrt{\frac{\bar{\gamma}_{\rho}+L_{\rho}}{\mu}}+\frac{m}{3}\sqrt{\frac{2(\bar{\gamma}_{\rho}+L_{\rho})}{\delta\gamma_{\min}-\mu}}\left(\frac{3}{m}+\frac{2(\bar{\gamma}_{\rho}+L_{\rho})}{\delta\gamma_{\min}}\right)\right)\chi\big{(}q(x^{l(k_{0})})-q(\bar{x})\big{)}$
$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ +\left(\frac{4m}{3}\sqrt{\frac{\bar{\gamma}_{\rho}+L_{\rho}}{\mu}}+\frac{m}{3}\sqrt{\frac{2(\bar{\gamma}_{\rho}+L_{\rho})}{\delta\gamma_{\min}-\mu}}+\sqrt{\frac{\delta\gamma_{\min}}{\delta\gamma_{\min}-\mu}}\right)\sqrt{\frac{2\big{(}q(x^{l(k_{0}-m-1)})-q(\bar{x})\big{)}}{\delta\gamma_{\min}}}$
satisfies $\alpha<\frac{1}{2}$, where $\rho>0$, $\bar{\gamma}_{\rho}>0$, $\mu>0$ are the constants defined in (4.5), (4.7), and (4.4), respectively, while $\delta>0$ and $\gamma_{\min}>0$ are the parameters from Algorithm 4.1.
Proof.
From the fact that $\{x^{k}\}_{\mathcal{K}}\to\bar{x}$, the decreasement of $\{q(x^{l(k)})\}$, we have $\|x^{k_{0}}-\bar{x}\|$ is sufficiently small by the sufficiently large $k_{0}\in\mathcal{K}$, and $q(x^{l(k)})-q(\bar{x})\to 0$ from Assumption 4.2. Then we have $\chi(q(x^{l(k)})-q(\bar{x}))\to 0$ from the continuity of $\chi$ and $\chi(0)=0$. Meanwhile, because $k_{0}-m-1$ is sufficiently large, we have $q(x^{l(k_{0}-m-1)})-q(\bar{x})$ is sufficiently small. Totally, we have $\alpha<\frac{1}{2}$. ∎
Taking the similar analysis with Lemma 3.8, we obtain the following result.
Lemma 4.6.
Let Assumption 3.2 and Assumption 4.2 hold, and let $\{x^{k}\}$ be any sequence generated by Algorithm 4.1. Suppose that $\{x^{k}\}_{\mathcal{K}}$ is a subsequence converging to some limit point $\bar{x}$, then
$\operatorname{dist}\big{(}0,\partial q(x^{k+1})\big{)}\leq\big{(}\bar{\gamma}_{\rho}+L_{\rho}\big{)}\|x^{k+1}-x^{k}\|$
holds for all sufficiently large $k\geq k_{0}$ such that $x^{k}\in B_{\rho}(\bar{x})$, where $\rho$ denotes the constant from (4.5), $\bar{\gamma}_{\rho}>0$ is the constant from (4.7), and $L_{\rho}>0$ is the Lipschitz constant of $\nabla f$ on $C_{\rho}$ from (4.6).
By employing the KL property, we next illustrate that the subsequential trajectory has a finite length.
Theorem 4.7.
Let Assumption 3.2 and Assumption 4.2 hold, and let $\{x^{k}\}$ be any sequence generated by Algorithm 4.1. Suppose that $\{x^{k}\}_{\mathcal{K}}$ is a subsequence converging to some limit point $\bar{x}$, and that $q$ has the KL property at $\bar{x}$. If $q(x^{l(k)})\neq q(\bar{x})$ for all $k\in\mathbb{N}$ and for any $v\geq k_{0}$ where $k_{0}$ is used to define $\alpha$ in Lemma 4.5, if $x^{k}\in B_{\alpha}(\bar{x})$ holds for all $k_{0}\leq k\leq v$, then we have
$\sum_{k=k_{0}}^{v}\|x^{l(k)}-x^{l(k)-1}\|\leq\frac{4m}{3}\left(\sqrt{\frac{2\big{(}q(x^{l(k_{0}-m-1)})-q(\bar{x})\big{)}}{\delta\gamma_{\min}}}+\frac{2(\bar{\gamma}_{\rho}+L_{\rho})}{\delta\gamma_{\min}}\chi\big{(}q(x^{l(k_{0})})-q(\bar{x})\big{)}\right).$
Proof.
Without loss of generality, we also assume that $k_{0}>\hat{k}+m$ (defined by (4.3)) and $k_{0}$ is sufficiently large to satisfy
$q(x^{l(k_{0})})<q(\bar{x})+\eta,$
and then
$q(\bar{x})<q(x^{l(k)})\leq q(x^{l(k_{0})})<q(\bar{x})+\eta\quad\forall k\geq k_{0}.$ (4.9)
Let $\chi:[0,\eta]\to[0,\infty)$ be the desingularization function which comes along with the validity of the KL property of $q$. Due to $\chi(0)=0$ and $\chi^{\prime}(t)>0$ for all $t\in(0,\eta)$, one obtains
$\chi\big{(}q(x^{l(k)})-q(\bar{x})\big{)}\geq 0\quad\forall k\geq k_{0}.$
For any $v\geq k_{0}$, if $x^{k}\in B_{\alpha}(\bar{x})$ holds for all $k_{0}\leq k\leq v$, one has $x^{l(k)}\in B_{\alpha+m\eta}(\bar{x})$ for all $k\geq k_{0}$ from (4.3), then KL property of $q$ at $\bar{x}$ yields that
$\chi^{\prime}\big{(}q(x^{l(k)})-q(\bar{x})\big{)}\operatorname{dist}\big{(}0,\partial q(x^{l(k)})\big{)}\geq 1\quad\forall k_{0}\leq k\leq v,$
which deduces from the concavity of $\chi$ that,
$\chi\big{(}q(x^{l(k)})-q(\bar{x})\big{)}-\chi\big{(}q(x^{l(k+m+1)})-q(\bar{x})\big{)}\geq\frac{q(x^{l(k)})-q(x^{l(k+m+1)})}{\operatorname{dist}(0,\partial q(x^{l(k)}))}\quad\forall k_{0}\leq k\leq v.$ (4.10)
By (4.2), we have
$\delta\frac{\gamma_{\min}}{2}\|x^{l(k+m+1)}-x^{l(k+m+1)-1}\|^{2}$ $\leq q(x^{l(l(k+m+1)-1)})-q(x^{l(k+m+1)})$ (4.11)
$\leq q(x^{l(k)})-q(x^{l(k+m+1)})\quad\forall k\in\mathbb{N}.$
Due to $\|x^{l(k)-1}-\bar{x}\|\leq\|x^{l(k)-1}-x^{l(k)}\|+\|x^{l(k)}-\bar{x}\|\leq\alpha+(m+1)\eta<\rho$ for all $k_{0}\leq k\leq v$ by (4.3), then Lemma 4.6 implies
$\operatorname{dist}\big{(}0,\partial q(x^{l(k)})\big{)}\leq(\bar{\gamma}_{\rho}+L_{\rho})\|x^{l(k)}-x^{l(k)-1}\|.$
Then (4.10) and (4.11) imply that
$\chi\big{(}q(x^{l(k)}-q(\bar{x})\big{)}-\chi\big{(}q(x^{l(k+m+1)})-q(\bar{x})\big{)}\geq\frac{\delta\frac{\gamma_{\min}}{2}\|x^{l(k+m+1)}-x^{l(k+m+1)-1}\|^{2}}{(\bar{\gamma}_{\rho}+L_{\rho})\|x^{l(k)}-x^{l(k)-1}\|}$ (4.12)
holds for all $k_{0}\leq k\leq v$ satisfying $x^{k}\in B_{\alpha}(\bar{x})$. Employing the inequality $\sqrt{ab}\leq\frac{1}{4}a+b$ for any $a\geq 0$ and $b\geq 0$ and setting $\tau:=\frac{2(\bar{\gamma}_{\rho}+L_{\rho})}{\delta\gamma_{\min}}$, we have
$\|x^{l(k+m+1)}-x^{l(k+m+1)-1}\|$
$\leq\sqrt{\tau\|x^{l(k)}-x^{l(k)-1}\|\big{(}\chi\big{(}q(x^{l(k)}-q(\bar{x})\big{)}-\chi\big{(}q(x^{l(k+m+1)})-q(\bar{x})\big{)}\big{)}}$
$\leq\frac{1}{4}\|x^{l(k)}-x^{l(k)-1}\|+\tau\big{(}\chi\big{(}q(x^{l(k)}-q(\bar{x})\big{)}-\chi\big{(}q(x^{l(k+m+1)})-q(\bar{x})\big{)}\big{)}$
for all $k_{0}\leq k\leq v$ satisfying $x^{k}\in B_{\alpha}(\bar{x})$. Then, summation yields
$\sum_{k=k_{0}+m+1}^{v+m+1}\|x^{l(k)}-x^{l(k)-1}\|=\sum_{k=k_{0}}^{v}\|x^{l(k+m+1)}-x^{l(k+m+1)-1}\|$
$\leq\frac{1}{4}\sum_{k=k_{0}}^{v}\|x^{l(k)}-x^{l(k)-1}\|+\tau\sum_{k=k_{0}}^{v}\chi\big{(}q(x^{l(k)}-q(\bar{x})\big{)}-\chi\big{(}q(x^{l(k+m+1)})-q(\bar{x})\big{)}$
$\leq\frac{1}{4}\sum_{k=k_{0}}^{v}\|x^{l(k)}-x^{l(k)-1}\|+\tau\sum_{k=k_{0}}^{k_{0}+m}\chi\big{(}q(x^{l(k)})-q(\bar{x})\big{)},$
equivalently,
$\frac{3}{4}\sum_{k=k_{0}+m+1}^{v+m+1}\|x^{l(k)}-x^{l(k)-1}\|\leq\frac{1}{4}\sum_{k=k_{0}}^{k_{0}+m}\|x^{l(k)}-x^{l(k)-1}\|+\tau\sum_{k=k_{0}}^{k_{0}+m}\chi\big{(}q(x^{l(k)})-q(\bar{x})\big{)},$
which implies from (4.2) and Proposition 4.3 (a) that
$\frac{3}{4}\sum_{k=k_{0}}^{v+m+1}\|x^{l(k)}-x^{l(k)-1}\|$ $\leq\sum_{k=k_{0}}^{k_{0}+m}\|x^{l(k)}-x^{l(k)-1}\|+\tau\sum_{k=k_{0}}^{k_{0}+m}\chi\big{(}q(x^{l(k)})-q(\bar{x})\big{)}$
$\leq\sum_{k=k_{0}}^{k_{0}+m}\sqrt{\frac{2\big{(}q(x^{l(l(k)-1)})-q(x^{l(k)})\big{)}}{\delta\gamma_{\min}}}+\tau\sum_{k=k_{0}}^{k_{0}+m}\chi\big{(}q(x^{l(k)})-q(\bar{x})\big{)}$
$\leq m\sqrt{\frac{2\big{(}q(x^{l(l(k_{0})-1)})-q(x^{l(k_{0}+m)})\big{)}}{\delta\gamma_{\min}}}+m\tau\chi\big{(}q(x^{l(k_{0})})-q(\bar{x})\big{)}$
$\leq m\sqrt{\frac{2\big{(}q(x^{l(k_{0}-m-1)})-q(x^{l(k_{0}+m)})\big{)}}{\delta\gamma_{\min}}}+m\tau\chi\big{(}q(x^{l(k_{0})})-q(\bar{x})\big{)}.$
Hence, we have
$\sum_{k=k_{0}}^{v}\|x^{l(k)}-x^{l(k)-1}\|\leq\frac{4m}{3}\left(\sqrt{\frac{2\big{(}q(x^{l(k_{0}-m-1)})-q(x^{l(k_{0}+m)})\big{)}}{\delta\gamma_{\min}}}+\tau\chi\big{(}q(x^{l(k_{0})})-q(\bar{x})\big{)}\right).$
Theorem 4.8.
Let Assumption 3.2 and Assumption 4.2,hold, and let $\{x^{k}\}$ be any sequence generated by Algorithm 4.1. Suppose that $\{x^{k}\}_{\mathcal{K}}$ is a subsequence converging to some limit point $\bar{x}$, and that $q$ has the KL property at $\bar{x}$, then the entire sequence $\{x^{k}\}$ converges to $\bar{x}$.
Proof.
By Proposition 4.3 (a), one knows that the whole sequence $\{q(x^{l(k)})\}$ is monotonically decreasing and convergent to $q(\bar{x})$ by assumption. This implies that $q(x^{l(k)})\geq q(\bar{x})$ for all $k\in\mathbb{N}$. If $q(x^{l(\bar{k})})=q(\bar{x})$ holds for some index $\bar{k}\in\mathbb{N}$, which, by monotonicity, implies that $q(x^{l(\bar{k}+1)})=q(\bar{x})$, then we claim that $x^{k+1}=x^{k}$ for all $k\geq\bar{k}+m_{\bar{k}}$ ($m_{\bar{k}}=m$ if $\bar{k}$ is sufficiently large) by [25, Lemma 3.2 (iv)]. Since, by assumption, the subsequence $\{x^{k}\}_{\mathcal{K}}$ converges to $\bar{x}$, this implies that $x^{k}=\bar{x}$ for all $k\in\mathbb{N}$ sufficiently large. In particular, we have the convergence of the entire (eventually constant) sequence $\{x^{k}\}$ to $\bar{x}$ in this situation.
It remains to consider the case where $q(x^{l(k)})>q(\bar{x})$ for all $k\in\mathbb{N}$. Recall the analysis in Theorem 4.7, let us assume again that $k_{0}\geq\hat{k}$ and $k_{0}\in K$ is sufficiently large to satisfy
$q(x^{l(k_{0})})<q(\bar{x})+\eta.$ (4.13)
We now claim that the following two statements hold for all $k\geq k_{0}$:
  1. (a)
    $x^{k}\in B_{\alpha}(\bar{x})$,
  2. (b)
    $\|x^{k_{0}}-\bar{x}\|+\sum_{i=k_{0}}^{k}\|x^{i+1}-x^{i}\|\leq\alpha$, which is equivalent to
    $\sum_{i=k_{0}}^{k}\|x^{i+1}-x^{i}\|$ (4.14)
    $\leq\left(\frac{8m(\bar{\gamma}_{\rho}+L_{\rho})}{3\delta\gamma_{\min}}\sqrt{\frac{\bar{\gamma}_{\rho}+L_{\rho}}{\mu}}+\frac{m}{3}\sqrt{\frac{2(\bar{\gamma}_{\rho}+L_{\rho})}{\delta\gamma_{\min}-\mu}}\left(\frac{3}{m}+\frac{2(\bar{\gamma}_{\rho}+L_{\rho})}{\delta\gamma_{\min}}\right)\right)\chi\big{(}q(x^{l(k_{0})})-q(\bar{x})\big{)}$
    $\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ +\left(\frac{4m}{3}\sqrt{\frac{\bar{\gamma}_{\rho}+L_{\rho}}{\mu}}+\frac{m}{3}\sqrt{\frac{2(\bar{\gamma}_{\rho}+L_{\rho})}{\delta\gamma_{\min}-\mu}}+\sqrt{\frac{\delta\gamma_{\min}}{\delta\gamma_{\min}-\mu}}\right)\sqrt{\frac{2\big{(}q(x^{l(k_{0}-m-1)})-q(\bar{x})\big{)}}{\delta\gamma_{\min}}},$
where $\alpha$ is defined in (4.8). We still verify these two statements jointly by induction. For $k=k_{0}$, statement (a) holds by the definition of $\alpha$ in (4.8). Meanwhile, due to $k_{0}\in\overline{K}$, (4.4) and (4.1) deduce that
$\frac{\delta\gamma_{\min}-\mu}{2}\|x^{k_{0}+1}-x^{k_{0}}\|^{2}$ $\leq q(x^{l(k_{0})})-q(x^{k_{0}+1})+q(x^{k_{0}+1})-q(x^{l(k_{0}+1)})$
$=q(x^{l(k_{0})})-q(x^{l(k_{0}+1)})\leq q(x^{l(k_{0})})-q(\bar{x})$
which says that (4.14) also holds for $k=k_{0}$. Suppose that both statements are valid for all $k\geq k_{0}$. Using the triangle inequality, the induction hypothesis, and the definition of $\alpha$, we have
$\|x^{k+1}-\bar{x}\|\leq\sum_{i=k_{0}}^{k}\|x^{i+1}-x^{i}\|+\|x^{k_{0}}-\bar{x}\|\leq\alpha,$
i.e., statement (a) holds for $k+1$ in place of $k$. Therefore, we have $x^{i}\in B_{\alpha}(\bar{x})$ for all $k_{0}\leq i\leq k+1$ by induction. The verification of the induction step for (4.14) is more involved. We next consider two cases whether or not $k\in K$ defined in (4.4) for all $k\geq k_{0}$.
Case 1: $k\in\overline{K}$. For all $k$ satisfying $k\geq k_{0}$ and $k\in\overline{K}$, then one has
$0\leq q(x^{l(k+1)})-q(x^{k+1})\leq\frac{\mu}{2}\|x^{k+1}-x^{k}\|^{2}.$ (4.15)
Since KL property holds at $\bar{x}$, then (4.10) with $m=0$ is also valid, i.e.,
$q(x^{l(i)})-q(x^{l(i+1)})\leq\left(\chi\big{(}q(x^{l(i)})-q(\bar{x})\big{)}-\chi\big{(}q(x^{l(i+1)})-q(\bar{x})\big{)}\right)(\bar{\gamma}_{\rho}+L_{\rho})\|x^{l(i)}-x^{l(i)-1}\|$ (4.16)
for all $k_{0}\leq i\leq k+1$ and $i\in\overline{K}$. (4.1) and (4.15) imply that
$\frac{\delta\gamma_{k}-\mu}{2}\|x^{i+1}-x^{i}\|^{2}\leq q(x^{l(i)})-q(x^{i+1})+q(x^{i+1})-q(x^{l(i+1)})=q(x^{l(i)})-q(x^{l(i+1)}).$ (4.17)
Define
$\Delta_{i,j}:=\chi\big{(}q(x^{l(i)})-q(\bar{x})\big{)}-\chi\big{(}q(x^{l(j)})-q(\bar{x})\big{)}$
for short. Then, (4.16) and (4.17) yield that
$\frac{\delta\gamma_{\min}-\mu}{2}\|x^{i+1}-x^{i}\|^{2}\leq(\bar{\gamma}_{\rho}+L_{\rho})\Delta_{i,i+1}\|x^{l(i)}-x^{l(i)-1}\|$
for all $k_{0}\leq i\leq k+1$ and $i\in\overline{K}$. It implies from the inequality $\sqrt{ab}\leq\frac{1}{4}a+b$ for any $a\geq 0$ and $b\geq 0$ again that
$\|x^{i+1}-x^{i}\|\leq\sqrt{\frac{2(\bar{\gamma}_{\rho}+L_{\rho})}{\delta\gamma_{\min}-\mu}}\sqrt{\Delta_{i,i+1}\|x^{l(i)}-x^{l(i)-1}\|}\leq\hat{\tau}\left(\Delta_{i,i+1}+\frac{1}{4}\|x^{l(i)}-x^{l(i)-1}\|\right)$
for all $k_{0}\leq i\leq k+1$ and $i\in\overline{K}$ with $\hat{\tau}:=\sqrt{\frac{2(\bar{\gamma}_{\rho}+L_{\rho})}{\delta\gamma_{\min}-\mu}}$. Recall that $x^{i}\in B_{\alpha}(\bar{x})$ for all $k_{0}\leq i\leq k+1$ by induction hypothesis and hence Theorem 4.7 holds with $v=k+1$. Summation implies that
$\sum^{k+1}_{k_{0}=i\in\overline{K}}\leq\hat{\tau}\sum_{k_{0}=i\in\overline{K}}^{k+1}\left(\Delta_{i,i+1}+\frac{1}{4}\|x^{l(i)}-x^{l(i)-1}\|\right)\leq\hat{\tau}\sum_{i=k_{0}}^{k+1}\left(\Delta_{i,i+1}+\frac{1}{4}\|x^{l(i)}-x^{l(i)-1}\|\right)$ (4.18)
$\leq\hat{\tau}\chi\big{(}q(x^{l(k_{0})})-q(\bar{x})\big{)}+\hat{\tau}\frac{1}{4}\sum_{i=k_{0}}^{k+1}\|x^{l(i)}-x^{l(i)-1}\|$
$\leq\hat{\tau}\chi\big{(}q(x^{l(k_{0})})-q(\bar{x})\big{)}$
$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ +\frac{\hat{\tau}m}{3}\left(\sqrt{\frac{2\big{(}q(x^{l(k_{0}-m-1)})-q(\bar{x})\big{)}}{\delta\gamma_{\min}}}+\frac{2(\bar{\gamma}_{\rho}+L_{\rho})}{\delta\gamma_{\min}}\chi\big{(}q(x^{l(k_{0})})-q(\bar{x})\big{)}\right)$
$=\frac{m}{3}\sqrt{\frac{2(\bar{\gamma}_{\rho}+L_{\rho})}{\delta\gamma_{\min}-\mu}}\left(\sqrt{\frac{2\big{(}q(x^{l(k_{0}-m-1)})-q(\bar{x})\big{)}}{\delta\gamma_{\min}}}+\left(\frac{3}{m}+\frac{2(\bar{\gamma}_{\rho}+L_{\rho})}{\delta\gamma_{\min}}\right)\chi\big{(}q(x^{l(k_{0})})-q(\bar{x})\big{)}\right).$
Case 2: $k\in K$. For sufficiently large $k_{0}$, from (4.3), one has
$\|x^{l(k+1)}-x^{k+1}\|\leq\sum_{i=k+1-m}^{k+1}\|x^{i+1}-x^{i}\|\leq m\eta,\quad\forall k\geq k_{0},$ (4.19)
which means that $\|x^{l(k+1)}-x^{k+1}\|\to 0$. From (4.1), one has
$\gamma_{k}(x^{k}-x^{k+1})+\nabla f(x^{k+1})-\nabla f(x^{k})\in\hat{\partial}q(x^{k+1})\quad\forall k\in\mathbb{N}.$
Meanwhile, for all $k\in K$ where $K$ is defined in (4.4), we know that $x^{l(k+1)}\neq x^{k+1}$. We obtain that $\|x^{l(k+1)}-x^{k+1}\|\to 0$ by Theorem 4.7. Hence, by the definition of (Fréchet) subdifferential, we have
$q(x^{k+1})-q(x^{l(k+1)})-$ (4.20)
$\gamma_{l(k+1)-1}(x^{l(k+1)-1}-x^{l(k+1)})+\nabla f(x^{l(k+1)})-\nabla f(x^{l(k+1)-1}),x^{k+1}-x^{l(k+1)}\right>$
$\geq o(\|x^{k+1}-x^{l(k+1)}\|)$
Recall again that $x^{i}\in B_{\alpha}(\bar{x})$ for all $k_{0}\leq i\leq k+1$, then from (4.19), we have
$\|x^{l(i+1)}-\bar{x}\|$ $\leq\|x^{l(i+1)}-x^{i+1}\|+\|x^{i+1}-x^{i}\|+\|x^{i}-\bar{x}\|$
$\leq(m+1)\eta+\alpha\leq\rho\quad\forall k_{0}\leq i\leq k+1,$
and consequently
$\|x^{l(i+1)-1}-\bar{x}\|\leq\|x^{l(i+1)-1}-x^{l(i+1)}\|+\|x^{l(i+1)}-\bar{x}\|\leq\eta+(m+1)\eta+\alpha<2\rho$
for all $k_{0}\leq i\leq k+1$. Therefore, for all $k_{0}\leq i\leq k+1$, we have $\gamma_{l(i+1)-1}\leq\bar{\gamma}_{\rho}$ from Lemma 4.4 that $\gamma_{l(i+1)-1}\leq\bar{\gamma}_{\rho}$, and $x^{l(i+1)-1},x^{l(i+1)}\in C_{\rho}$, then we have
$\|\nabla f(x^{l(i+1)})-\nabla f(x^{l(i+1)-1})\|\leq L_{\rho}\|x^{l(i+1)}-x^{l(i+1)-1}\|.$
Therefore, (4.20) implies that
$q(x^{i+1})-q(x^{l(i+1)})+(\bar{\gamma}_{\rho}+L_{\rho})\|x^{l(i+1)-1}-x^{l(i+1)}\|\|x^{i+1}-x^{l(i+1)}\|\geq o(\|x^{i+1}-x^{l(i+1)}\|)$
for all $k_{0}\leq i\leq k+1$ and $i\in K$. Employing the inequality $a^{2}+b^{2}\geq 2ab$ for all $a,b\in\mathbb{R}$ as well as $\|x^{i+1}-x^{l(i+1)}\|\to 0$ yields that
$q(x^{i+1})-q(x^{l(i+1)})+\frac{\bar{\gamma}_{\rho}+L_{\rho}}{2}\left(\|x^{l(i+1)-1}-x^{l(i+1)}\|^{2}+\|x^{i+1}-x^{l(i+1)}\|^{2}\right)$
$\geq\frac{\bar{\gamma}_{\rho}+L_{\rho}}{2}\|x^{i+1}-x^{l(i+1)}\|^{2}\quad\forall k_{0}\leq i\leq k+1\ \text{and}\ i\in K,$
which implies that
$\frac{\bar{\gamma}_{\rho}+L_{\rho}}{2}\|x^{l(i+1)-1}-x^{l(i+1)}\|^{2}\geq q(x^{l(i+1)})-q(x^{i+1})$ (4.21)
for all $k_{0}\leq i\leq k+1$ and $i\in K$. By (4.21) and (4.4), we have
$\frac{\mu}{2}\|x^{i+1}-x^{i}\|^{2}<q(x^{l(i+1)})-q(x^{i+1})\leq\frac{\bar{\gamma}_{\rho}+L_{\rho}}{2}\|x^{l(i+1)-1}-x^{l(i+1)}\|^{2}$
for all $k_{0}\leq i\leq k+1$ and $i\in K$, which definitely says that
$\|x^{i+1}-x^{i}\|\leq\sqrt{\frac{\bar{\gamma}_{\rho}+L_{\rho}}{\mu}}\|x^{l(i+1)-1}-x^{l(i+1)}\|\quad\forall k_{0}\leq i\leq k+1\ \text{and}\ i\in K.$
We know that the induction hypothesis yields Theorem 4.7 holds automatically for $v=k+1$, hence, we have
$\sum_{k_{0}=i\in K}^{k+1}\|x^{i+1}-x^{i}\|\leq\sqrt{\frac{\bar{\gamma}_{\rho}+L_{\rho}}{\mu}}\sum_{k_{0}=i\in K}^{k+1}\|x^{l(i+1)-1}-x^{l(i+1)}\|$ (4.22)
$\leq\sqrt{\frac{\bar{\gamma}_{\rho}+L_{\rho}}{\mu}}\sum_{i=k_{0}}^{k+1}\|x^{l(i+1)-1}-x^{l(i+1)}\|$
$<\frac{4m}{3}\sqrt{\frac{\bar{\gamma}_{\rho}+L_{\rho}}{\mu}}\left(\sqrt{\frac{2\big{(}q(x^{l(k_{0}-m-1)})-q(\bar{x})\big{)}}{\delta\gamma_{\min}}}+\frac{2(\bar{\gamma}_{\rho}+L_{\rho})}{\delta\gamma_{\min}}\chi\big{(}q(x^{l(k_{0})})-q(\bar{x})\big{)}\right).$
By (4.18) and (4.22), we have
$\sum_{i=k_{0}}^{k+1}\|x^{i+1}-x^{i}\|\leq\sum_{k_{0}=i\in K}^{k+1}\|x^{i+1}-x^{i}\|+\sum_{k_{0}=i\in\overline{K}}^{k+1}\|x^{i+1}-x^{i}\|$
$\leq\left(\frac{8m(\bar{\gamma}_{\rho}+L_{\rho})}{3\delta\gamma_{\min}}\sqrt{\frac{\bar{\gamma}_{\rho}+L_{\rho}}{\mu}}+\frac{m}{3}\sqrt{\frac{2(\bar{\gamma}_{\rho}+L_{\rho})}{\delta\gamma_{\min}-\mu}}\left(\frac{3}{m}+\frac{2(\bar{\gamma}_{\rho}+L_{\rho})}{\delta\gamma_{\min}}\right)\right)\chi\big{(}q(x^{l(k_{0})})-q(\bar{x})\big{)}$
$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ +\left(\frac{4m}{3}\sqrt{\frac{\bar{\gamma}_{\rho}+L_{\rho}}{\mu}}+\frac{m}{3}\sqrt{\frac{2(\bar{\gamma}_{\rho}+L_{\rho})}{\delta\gamma_{\min}-\mu}}\right)\sqrt{\frac{2\big{(}q(x^{l(k_{0}-m-1)})-q(\bar{x})\big{)}}{\delta\gamma_{\min}}}$
$\leq\left(\frac{8m(\bar{\gamma}_{\rho}+L_{\rho})}{3\delta\gamma_{\min}}\sqrt{\frac{\bar{\gamma}_{\rho}+L_{\rho}}{\mu}}+\frac{m}{3}\sqrt{\frac{2(\bar{\gamma}_{\rho}+L_{\rho})}{\delta\gamma_{\min}-\mu}}\left(\frac{3}{m}+\frac{2(\bar{\gamma}_{\rho}+L_{\rho})}{\delta\gamma_{\min}}\right)\right)\chi\big{(}q(x^{l(k_{0})})-q(\bar{x})\big{)}$
$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ +\left(\frac{4m}{3}\sqrt{\frac{\bar{\gamma}_{\rho}+L_{\rho}}{\mu}}+\frac{m}{3}\sqrt{\frac{2(\bar{\gamma}_{\rho}+L_{\rho})}{\delta\gamma_{\min}-\mu}}+\sqrt{\frac{\delta\gamma_{\min}}{\delta\gamma_{\min}-\mu}}\right)\sqrt{\frac{2\big{(}q(x^{l(k_{0}-m-1)})-q(\bar{x})\big{)}}{\delta\gamma_{\min}}}$
Hence, statement (b) holds for $k+1$, and this completes the induction.
In particular, it follows from (a) that $x^{k}\in B_{\alpha}(\bar{x})$ for all $k\geq k_{0}$. Taking $k\to\infty$ in (4.14) therefore shows that $\{x^{k}\}$ is a Cauthy sequence and, thus, convergent. Since we already know that $\bar{x}$ is an accumulation point, it follows that the entire sequence $\{x^{k}\}$ is convergent to $\bar{x}$. ∎
Note that if $\overline{K}=\mathbb{N}$ or $m=0$, then we have $q(x^{k})>q(x^{k+1})$ for all $k\in\mathbb{N}$, i.e., $\{q(x^{k})\}$ is decreasing, then Algorithm 4.1 degenerates into [15, Algorithm 3.1], where the specific results about the convergence and convergence rate of the whole sequence were proposed. In the following, we illustrate the more general convergence rate of the whole sequence where $\overline{K}\neq\mathbb{N}$.
Theorem 4.9.
Let Assumption 3.2 and Assumption 4.2 hold, and let $\{x^{k}\}$ be any sequence generated by Algorithm 4.1. Suppose that $\{x^{k}\}_{\mathcal{K}}$ is a subsequence converging to some limit point $\bar{x}$, and that $q$ has the KL property at $\bar{x}$. Then the entire sequence $\{x^{k}\}$ converges to $\bar{x}$, and if the corresponding desingularization function has the form $\chi(t)=ct^{\theta}$ for some $c>0$ and $\theta\in(0,1]$, then the following statements hold:
  1. (a)
    if $\theta=1$, then the sequences $\{q(x^{l(k)})\}$ and $\{x^{k}\}$ converge in a finite numer of steps to $q(\bar{x})$ and $\bar{x}$, respectively.
  2. (b)
    if $\theta\in[\frac{1}{2},1)$, then the sequence $\{q(x^{l(k)})\}$ converges Q-linearly to $q(\bar{x})$, and the sequence $\{x^{k}\}$ converges R-linearly to $\bar{x}$.
  3. (c)
    if $\theta\in(0,\frac{1}{2})$, the there exist some positive constants $\eta_{1}$ and $\eta_{2}$ such that
    $q(x^{l(k)})-q(\bar{x})$ $\leq\eta_{1}k^{-\frac{1}{1-2\theta}}$
    $\|x^{k}-\bar{x}\|$ $\leq\eta_{2}k^{-\frac{\theta}{1-2\theta}}$
    for sufficiently large $k$.
Proof.
Without loss of generality, we assume that $q(x^{l(k)})>q(\bar{x})$ for all $k\in\mathbb{N}$. Therefore, for all $k\geq k_{0}$ which is defined by Lemma 4.5, one has (4.9) holds, and then KL property of $q$ at $\bar{x}$ and Lemma 4.6 imply that
11 11 $\leq\chi^{\prime}\big{(}q(x^{l(k)})-q(\bar{x})\big{)}\operatorname{dist}\big{(}0,\partial q(x^{l(k)})\big{)}$ (4.23)
$\leq c\theta\big{(}q(x^{l(k)})-q(\bar{x})\big{)}^{\theta-1}(\bar{\gamma}_{\rho}+L_{\rho})\|x^{l(k)}-x^{l(k)-1}\|\quad\forall k\geq k_{0}.$
(4.11) and (4.23) imply that
$q(x^{l(k+1)})-q(x^{l(k)})$ $\leq-\delta\frac{\gamma_{\min}}{2}\|x^{l(k+1)}-x^{l(k+1)-1}\|^{2}$
$\leq-\delta\frac{\gamma_{\min}}{2(c\theta)^{2}(\bar{\gamma}_{\rho}+L_{\rho})^{2}}\big{(}q(x^{l(k+1)})-q(\bar{x})\big{)}^{2(1-\theta)}$
$:=-\xi\big{(}q(x^{l(k+1)})-q(\bar{x})\big{)}^{2(1-\theta)}\quad\forall k\geq k_{0}$
with $\xi:=\delta\frac{\gamma_{\min}}{2(c\theta)^{2}(\bar{\gamma}_{\rho}+L_{\rho})^{2}}$. Hence, we have
$\big{(}q(x^{l(k+1)})-q(\bar{x})\big{)}^{2(1-\theta)}\leq\frac{1}{\xi}\big{(}q(x^{l(k)})-q(x^{l(k+1)})\big{)}\quad\forall k\geq k_{0}.$ (4.24)
Since $\{q(x^{l(k)})\}$ is decreasing, then statements (a), (b), and (c) regarding the sequence $\{q(x^{l(k)})\}$ follow from [14, Lemma 2.14]. It remains to verify the convergence rate with respect to the sequence $\{x^{k}\}$. We now consider the different cases of $\theta$.
  • $\theta=1$: for all $k\geq k_{0}$, (4.24) implies that $q(x^{l(k+1)})-q(x^{l(k)})\leq-\xi$, which yields that $\{q(x^{l(k)})\}$ converges to $q(\bar{x})$ in finite steps, then one has $\{x^{k}\}$ also converges to $\bar{x}$ in finite steps.
  • $\theta\in(0,1/2)$: for all $k\geq k_{0}$ and $s>k$, (4.14) implies that
    $\|x^{k}-x^{s}\|$ $\leq\sum_{i=k}^{s}\|x^{i+1}-x^{i}\|\leq a\sqrt{q(x^{l(k-m-1)})-q(\bar{x})}+bc\big{(}q(x^{l(k)})-q(\bar{x})\big{)}^{\theta}$
    $\leq(a+bc)\big{(}q(x^{l(k-m-1)})-q(\bar{x})\big{)}^{\theta},$
    where $a=\left(\frac{4m}{3}\sqrt{\frac{\bar{\gamma}_{\rho}+L_{\rho}}{\mu}}+\frac{m}{3}\sqrt{\frac{2(\bar{\gamma}_{\rho}+L_{\rho})}{\delta\gamma_{\min}-\mu}}+\sqrt{\frac{\delta\gamma_{\min}}{\delta\gamma_{\min}-\mu}}\right)\sqrt{\frac{2}{\delta\gamma_{\min}}}$, $b=\frac{8m(\bar{\gamma}_{\rho}+L_{\rho})}{3\delta\gamma_{\min}}\sqrt{\frac{\bar{\gamma}_{\rho}+L_{\rho}}{\mu}}+\frac{m}{3}\sqrt{\frac{2(\bar{\gamma}_{\rho}+L_{\rho})}{\delta\gamma_{\min}-\mu}}\left(\frac{3}{m}+\frac{2(\bar{\gamma}_{\rho}+L_{\rho})}{\delta\gamma_{\min}}\right)$.
    Note that there exists some constant $\eta_{1}$ such that
    $q(x^{l(k)})-q(\bar{x})\leq\eta_{1}k^{\frac{1}{2\theta-1}}\quad\forall\ \text{sufficiently large}\ k,$
    then
    $\|x^{k}-x^{s}\|\leq(a+b)\eta_{1}^{\theta}k^{\frac{\theta}{2\theta-1}}:=\eta_{2}k^{\frac{\theta}{2\theta-1}}\quad\forall\ \text{sufficiently large}\ k,$
    where $\eta_{2}:=(a+b)\eta_{1}^{\theta}$, taking $s\to\infty$ yields the desired result.
  • $\theta\in[1/2,1)$: for all $k\geq k_{0}$ and $s>k$, (4.14) implies that
    $\|x^{k}-x^{s}\|$ $\leq\sum_{i=k}^{s}\|x^{i+1}-x^{i}\|\leq a\sqrt{q(x^{l(k-m-1)})-q(\bar{x})}+bc\big{(}q(x^{l(k)})-q(\bar{x})\big{)}^{\theta}$
    $\leq(a+bc)\big{(}q(x^{l(k-m-1)})-q(\bar{x})\big{)}^{\frac{1}{2}}.$
    Note that $\{q(x^{l(k)})\}$ Q-linearly converges to $q(\bar{x})$, then $\{x^{k}\}$ R-linearly converges to $\bar{x}$ (by taking $s\to\infty$).
This completes the proof. ∎
5 Conclusion
This manuscript presented the convergence and rate-of-convergence results of two types of nonmonotine proximal gradient methods, i.e., combined with average line search and max line search, respectively. These results were established under the KL property, requiring only that the smooth part of the objective function has locally (rather than globally) Lipschitz continuous gradients. Note that the partitioning of the index set is a useful tool for the convergence of NPG. At least for NPG with average line search, the convergence theory is independent on the specific choice of partitioning. Additionally, our results in Section 4 address the question raised in [17, Final Remarks], demonstrating that the technique of proof there can indeed be extended to NPG with max line search.

### v3

@@
-Advances in Nonmonotone Proximal Gradient Methods merely with Local Lipschitz Assumptions in the Presense of Kurdyka–Łojasiewicz Property: A Study of Average and Max Line Search
+Convergence analysis of nonmonotone proximal gradient methods under local Lipschitz continuity and Kurdyka–Łojasiewicz property
 Abstract.
-The proximal gradient method is a standard approach to solve the composite minimization problems where the objective function is the sum of a continuously differentiable function and a lower semicontinuous, extended-valued function.
-For both monotone and nonmonotone proximal gradient methods, the convergence theory has traditionally replied heavily on the assumption of global Lipschitz continuity.
-Recent works have shown that the monotone proximal gradient method, even when the local Lipschitz continuity (rather than global) is assumed, converges to the stationarity globally in the presence of Kurdyka–Łojasiewicz Property.
-However, how to extend these results from monotone proximal gradient method to nonmonotone proximal gradient method (NPG) remains an open question.
-In this manuscript, we consider two types of NPG: those combined with average line search and max line search, respectively.
-By partitioning of indices into two subsets, one of them aims to achieve a decrease in the functional sequence, we establish the global convergence and rate-of-convergence (same as the monotone version) results under the KL property, merely requiring the local Lipschitz assumption, and without an a priori knowledge of the iterative sequence being bounded.
-When our work is almost done, we noticed that [17] presented the analogous results for the NPG with average line search, whose partitioning of index set is totally different with ours.
-Drawing upon the findings in this manuscript and [17], we confidently conclude that the convergence theory of NPG is independent on the specific partitioning of the index set.
-Keywords. Composite problems $\cdot$ Nonmonotone proximal gradient methods $\cdot$ Kurdyka-Łojasiewicz property $\cdot$ Local Lipschitz continuity
+The proximal gradient method is a standard approach for solving composite minimization problems in which the objective function is the sum of a continuously differentiable function and a lower semicontinuous, extended-valued function.
+The traditional convergence theory for both monotone and nonmonotone variants replies heavily on the assumption of global Lipschitz continuity of the gradient of the smooth part of the objective function.
+Recent work has shown that monotone proximal gradient methods converge globally only when the local (rather than global) Lipschitz continuity is assumed, provided that the Kurdyka–Łojasiewicz (KL) property holds.
+However, these results have not been extended to nonmonotone proximal gradient (NPG) methods.
+In this manuscript, we consider two types of NPG methods: those combined with the average line search and the max line search, respectively.
+By partitioning indices into two subsets, one of which aims to achieve a sufficient decrease in the functional sequence, we establish global convergence and rate-of-convergence results using the local Lipschitz continuity and the KL property, without requiring boundedness of the iterates.
+While finalizing this work, we noticed that [18] presented analogous results for the NPG method with average line search, but with a different partitioning strategy.
+Together, we confidently conclude that the convergence theory of the NPG method is independent on index partitioning choices.
+Keywords. Composite problems $\cdot$ Nonmonotone proximal gradient methods $\cdot$ Average and max line searches $\cdot$ Kurdyka-Łojasiewicz property $\cdot$ Local Lipschitz continuity
 AMS subject classifications. 49J52, 90C26, 90C30
 1 Introduction
-Let us consider
+Let us consider the following optimization problem
 $\min_{x}\ q(x):=f(x)+g(x)\qquad\text{s.t.}\quad x\in\mathbb{X},$ (Q)
 where $\mathbb{X}$ is an Euclidean space, i.e., real and finite-dimensional Hilbert space, $f:\mathbb{X}\to\mathbb{R}$ is continuously differentiable, $g:\mathbb{X}\to\overline{\mathbb{R}}$ is assumed to be merely lower semicontinuous.
-We claim that the unconstrained composite optimization problem (Q) is in a totally nonconvex setting.
-This type of optimization problems have a wide range of applications in practice, like machine learning, image processing, and data science [4, 11, 8].
-In this manuscript, we are interested in two classes of nonmonotone proximal gradient method to solve (Q), where the nonmonotone line search technique based on monitoring the weighted average of the objective value at all iterates by Zhang and Hager [28], and the maximum objective value attained from the latest iterates by Grippo et al.
-[13], respectively.
-More precisely, some iterate $x^{k+1}\in\operatorname{dom}q$ is acceptable provided that the corresponding objective $q(x^{k+1})$ is less than the merit
-$\Phi_{k}:=(1-p)\Phi_{k-1}+pq(x^{k})\qquad\text{and}\qquad q(x^{l(k)}):=\max_{j=0,\ldots,\min\{m,k\}}q(x^{k-j})$
-for the average and max taste, respectively.
-Note that $m\in\mathbb{N}$ and $p\in(0,1]$ (usually) are constants in order to control the level of nonmonotonicity, i.e., for the smaller $p$, or the larger $m$, the corresponding metric function can attain larger values, hence imposing a weaker condition for acceptance of $x^{k+1}$ compared to the monotone proximal gradient methods [12].
-The values of $p=1$ and $m=0$ lead to the monotone behavior, i.e., the nonmotone proximal gradient method degenerates into the monotone one, please see [18, Algorithm 3.1] and [15, Algorithm 3.1].
-The convergence theory associated with (monotone and nonmonotone) proximal gradient methods mostly require the derivative of the smooth part of the objective function to be global Lipschitz continuous, which, however, is quite restrictive assumption in practical scenarios.
-[15, Examples 3.6 and 3.7], as two classical examples, revealed that the global Lipschitz assumption on the gradient of $f$ is typically violated, whereas the local Lipschitz condition is often satisfied.
-[18] readdressed the classical monotone proximal gradient method [18, Algorithm 3.1] and NPG with max line search [18, Algorithm 4.1], and provided (subsequential) convergence results only with local Lipschitz assumption (not global any more).
-As for the convergence of the whole sequence generated by the proximal gradient method with local Lipschitz assumption, [15] is the first one to reveal this mystery for monotone proximal gradient in the presence of the Kurdyka-Łojasiewicz property of the objective function.
-[12] considered an adaptive nonmonotone proximal gradient scheme based on an averaged metric function and established (subsequential) convergence results with local Lipschitz assumption of the gradient of $f$, where the global worst-case rates for the iterates and a (subsequential) stationarity measure were also derived.
-Note that the numerical behaviours and competitiveness has already been investigated in many literature before, for example see [26, 12, 25, 20, 21, 27], this manuscript focuses more on theoretical aspects of the method (without numerical experiments), showing qualitative and even quantitative properties of the NPG in a more general framework which allows for applications e.g.
-in the context of multiplier-penalty methods [16].
-More precisely, we, in this manuscript, under the same assumptions, consider the convergence theory of the entire sequence generated by NPG with both average and max backtracking line search when KL property being employed.
-Note that the corresponding convergence results about NPG with max line search has been analysed in [25], where an auxiliary sequence is necessarily prerequisite to be convergent, which, however, is not needed at all in our technique, moreover, where the boundedness of subdifferential of the objective function is needed for all iterates, which, to the best of our knowledge, fails to be satisfied under the local Lipschitz continuous assumption.
-However, in this manuscript, no any prerequisite is assumed and the corresponding convergence and rate-of-convergence results obtained in [25] can be also achieved merely by requiring the local Lipschitz continuity of $\nabla f$.
-The primary challenge posed by NPG, in fact, is that the generated functional sequence may not be decreasing, to date, whereas the (sufficient) descent property seems to be necessary for the technical proof when using KL property [2, 3, 7, 8].
-To the best of our knowledge, nearby all existing studies addressing this issue reply on the partition of the index sets [25, 17, 26].
-Specifically, these works divided the indices into two subsets: one subset ensures a decrease in the functional sequence, while the complement accounts for the remaining cases.
-The recent paper [17] considered the convergence results of the NPG with average line search under the KL property, merely requiring the local Lipschitz continuity of $\nabla f$, where the index sets are defined as
-$K_{1}:=\{k\in\mathbb{N}\,|\,q(x^{k})\leq\Phi_{k+m}\},\quad K_{2}:=\{k\in\mathbb{N}\,|\,q(x^{k})>\Phi_{k+m}\},$
-where $m$ is some predetermined integer relying on $p_{\min}$ to facilitate the convergence analysis.
-This partitioning originates from [26] and the subsequent convergence analysis heavily depend on the techniques therein, where the key component is exploiting the quantitative relationship between $\sum_{i=k}^{k+m-1}\sqrt{\Phi_{i-1}-\Phi_{i}}$ and $\sqrt{\Phi_{i-1}-\Phi_{i}}$, see [17, Lemma 4.4] and [26, Theorem 2.1].
+Note that (Q) is totally nonconvex, which has a wide range of applications in practice, like machine learning, image processing, and data science [4, 12, 8].
+In this manuscript, we are interested in two types of nonmonotone proximal gradient methods for solving (Q), where the nonmonotone line search techniques are based on different nonmonotone criteria.
+Specifically, we consider the approach of Zhang and Hager [29], which monitors the weighted average of the objective values over the past iterates, and the approach of Grippo et al.
+[14], which uses the maximum objective value attained among the most recent iterates.
+In both cases, a candidate iterate $x^{k+1}\in\operatorname{dom}q$ is acceptable if the corresponding objective $q(x^{k+1})$ is less than a designated merit function, defined as
+[…]
+where $m\in\mathbb{N}$ and $p\in(0,1]$ are constants, $l(k)\in\{k-m,\ldots,k\}$.
+Note that $p$ and $m$ aim to control the level of nonmonotonicity.
+For the smaller $p$, or the larger $m$, the corresponding metric function attains larger values, hence imposing a weaker condition for the acceptance of $x^{k+1}$ compared to the monotone proximal gradient methods [13].
+The values of $p=1$ and $m=0$ lead to the monotone behavior, i.e., the nonmotone proximal gradient method degenerates into the monotone one, please see [19, Algorithm 3.1] and [16, Algorithm 3.1].
+The convergence theory associated with (monotone and nonmonotone) proximal gradient methods primarily requires the derivative of the smooth part of the objective function to be globally Lipschitz continuous.
+However, the requirement is quite restrictive in practical scenarios.
+[16, Examples 3.6 and 3.7], as two classical examples, revealed that the global Lipschitz assumption on $\nabla f$ is typically violated, whereas the local Lipschitz condition is often satisfied.
+[19] readdressed the classical monotone proximal gradient method [19, Algorithm 3.1] and the NPG method with a max line search [19, Algorithm 4.1], and provided subsequential convergence results based solely on the local (not global any more) Lipschitz assumption.
+Regarding the convergence of the entire sequence generated by proximal gradient methods using the local Lipschitz assumption, [16, 9] revealed this mystery for monotone proximal gradient methods in the presence of the Kurdyka-Łojasiewicz (KL) property of the objective function.
+[13] considered an adaptive nonmonotone proximal gradient scheme based on an averaged metric function and established the subsequential convergence results with the local Lipschitz assumption of $\nabla f$, where the global worst-case rates for the iterates and the subsequential stationarity were also derived.
+Note that numerical behaviours and competitiveness have already been investigated in many literature, for example see [27, 13, 26, 21, 22, 28].
+In this manuscript, we emphasize theoretical aspects of the method (without numerical experiments), showing qualitative and even quantitative properties of NPG methods with a more general framework, facilitating various applications, such as in the context of multiplier-penalty methods [17].
+More precisely, we, in this manuscript, consider the convergence theory of the entire sequence generated by the NPG methods with both average and max backtracking line search techniques under the same assumptions when the KL property is employed.
+The corresponding convergence results for the NPG method with max line search have been analyzed in [26], in which the presence of an auxiliary sequence is a necessary prerequisite for convergence.
+However, such requirement is not required in our technique.
+One the other hand, the boundedness of subdifferential of the objective function is needed in [26] for all iterates, which, to the best of our knowledge, does not hold under the local Lipschitz continuity.
+In this manuscript, note that no any prerequisite is assumed and the corresponding convergence and rate-of-convergence results obtained in [26] can be also achieved by merely requiring the local Lipschitz continuity of $\nabla f$.
+The primary challenge posed by the NPG method, in fact, is that the generated functional sequence may not exhibit a sufficient decrease.
+To date, the sufficient descent property appears to be necessary for the technical proof when the KL property is employed [2, 3, 7, 8].
+To the best of our knowledge, nearby all existing studies addressing this issue reply on partitioning the index sets [26, 18, 27].
+Specifically, these works divide the indices into two subsets: one subset ensures a sufficient decrease in the functional sequence, while the complement accounts for the remaining case.
+A recent paper [18] considered the convergence results of the NPG method with average line search under the KL property, merely requiring the local Lipschitz continuity of $\nabla f$, where the index sets are defined as follows:
+$K_{1}:=\{k\in\mathbb{N}\,|\,q(x^{k})\leq\Phi_{k+\bar{m}}\},\quad K_{2}:=\{k\in\mathbb{N}\,|\,q(x^{k})>\Phi_{k+\bar{m}}\},$
+where $\bar{m}$ is a predetermined integer that relies on $p_{\min}$ to facilitate the convergence analysis.
+This partitioning originates from [27] and the subsequent convergence analysis heavily depends on the techniques presented therein, where the key component is exploiting the quantitative relationship between $\sum_{i=k}^{k+\bar{m}-1}\sqrt{\Phi_{i-1}-\Phi_{i}}$ and $\sqrt{\Phi_{i-1}-\Phi_{i}}$, see [18, Lemma 4.4] and [27, Theorem 2.1].
 To be honest, we and [18] happened to consider the same topic almost simultaneously.
-In this manuscript, for the NPG with average line search, to the best of our knowledge, we are the first one to propose a different partitioning of the index sets from [26]:
+In this manuscript, for the NPG method with average line search, to the best of our knowledge, we are the first to propose a different partitioning of the index sets from [27, 18], as follows:
 $S:=\{k\in\mathbb{N}\,|\,q(x^{k})-\Phi_{k+1}\leq\frac{\mu}{2}\|x^{k+1}-x^{k}\|^{2}\},\quad\overline{S}:=\mathbb{N}\setminus S,$
-where $\mu\in(0,\frac{1}{2}\delta{p_{\min}\gamma_{\min}}]$ is a constant, and $p_{\min}$ and $\gamma_{\min}$ are parameters derived from the algorithm.
-Actually, different partitioning of the index sets necessitates the different techniques for proof.
-However, our approach yields more natural and concise proofs compared to [17], as it eliminates the need for the mentioned-above quantitative relationship ([17, Lemma 4.4] and [26, Theorem 2.1]).
-Our analysis reproduces the corresponding convergence and rate-of-convergence presented in [17].
-Furthermore, [17] leaves the theoretical analysis of NPG methods with max line search for future work, we provide a comprehensive analysis in this manuscript.
-The decision to present this part of work (NPG with average line search), despite its similarities to [17], is also justified by a noteworthy finding: both works demonstrate that the convergence results of nonmonotone proximal gradient methods are independent on the specific partitioning of the index sets.
-This insight contributes a significant advancement to the understanding of this method.
+where $\mu\in(0,\frac{1}{2}\delta{p_{\min}\gamma_{\min}}]$ is a constant, and $\delta$, $p_{\min}$, and $\gamma_{\min}$ are parameters derived from the algorithm.
+Actually, different partitionings of the index sets necessitate different techniques for proofs.
+However, our approach yields more natural and concise proofs than those in [18], because it eliminates the need for the mentioned-above quantitative relationship ([18, Lemma 4.4] and [27, Theorem 2.1]).
+Furthermore, [18] presents the theoretical analysis of the NPG method with a max line search for future work, we conduct the comprehensive analysis in this manuscript.
+The decision to present this part of work (the NPG method with average line search), despite its similarities to [18], is also justified by a noteworthy finding: both works demonstrate that the convergence results of NPG methods are independent on the specific partitioning of the index sets.
+This insight significantly advances the understanding of the methods.
 This manuscript is organized as follows: We first recall some background knowledge in Section 2.
-We then present the basic properties and convergence results of the NPG with average line search and with max line search in Section 3 and Section 4, respectively.
+We then present the basic properties and convergence results of NPG methods with average line search and max line search in Section 3 and Section 4, respectively.
 Section 5 concludes this manuscript.
@@
 $\operatorname{dist}(x,A):=\inf\{\left\|y-x\right\|\,|\,y\in A\}$
-denotes the distance of the point $x$ to the set $A$ with $\operatorname{dist}(x,\emptyset):=\infty$.
-For given $\varepsilon>0$, $B_{\varepsilon}(x):=\{y\in\mathbb{X}\,|\,\left\|y-x\right\|\leq\varepsilon\}$ denotes the closed $\varepsilon$-ball around $x$.
-The continuous linear operator $f^{\prime}(x)\colon\mathbb{X}\to\mathbb{R}$ denotes the derivative of the continuously differentiable function $f\colon\mathbb{X}\to\mathbb{R}$ at $x\in\mathbb{X}$, and we will make use of $\nabla f(x):=f^{\prime}(x)^{*}1$ where $f^{\prime}(x)^{*}\colon\mathbb{R}\to\mathbb{X}$ is the adjoint of $f^{\prime}(x)$.
-This way, $\nabla f$ is a mapping from $\mathbb{X}$ to $\mathbb{X}$.
-We further say that a sequence $\{x^{k}\}\subset\mathbb{X}$ converges Q-linearly to $x^{*}\in\mathbb{X}$ if there is a constant $c\in(0,1)$ such that the inequality
+as the distance of the point $x$ to the set $A$ with $\operatorname{dist}(x,\emptyset):=\infty$.
+For any given $\varepsilon>0$, denote $B_{\varepsilon}(x):=\{y\in\mathbb{X}\,|\,\left\|y-x\right\|\leq\varepsilon\}$ as the closed $\varepsilon$-ball centered at $x$.
+Let the function $f\colon\mathbb{X}\to\mathbb{R}$ be continuously differentiable, the continuous linear operator $f^{\prime}(x)\colon\mathbb{X}\to\mathbb{R}$ denotes its derivative at $x\in\mathbb{X}$, and we use $\nabla f(x):=f^{\prime}(x)^{*}1$ where $f^{\prime}(x)^{*}\colon\mathbb{R}\to\mathbb{X}$ is the adjoint of $f^{\prime}(x)$.
+So, $\nabla f$ is a mapping from $\mathbb{X}$ to $\mathbb{X}$.
+A sequence $\{x^{k}\}\subset\mathbb{X}$ is said to converge Q-linearly to $x^{*}\in\mathbb{X}$ if there is a constant $c\in(0,1)$ such that the inequality
 $\|x^{k+1}-x^{*}\|\leq c\|x^{k}-x^{*}\|$
-holds for all sufficiently large $k\in\mathbb{N}$. Furthermore, $\{x^{k}\}$ is said to converge R-linearly to $x^{*}$ if we have
+holds for all sufficiently large $k\in\mathbb{N}$. Furthermore, $\{x^{k}\}$ converges R-linearly to $x^{*}$ if and only if we have
 $\limsup_{k\to\infty}\|x^{k}-x^{*}\|^{1/k}<1.$
-Note that this R-linear convergence holds if there exist constants $\omega>0$ and $\mu\in(0,1)$ such that $\|x^{k}-x^{*}\|\leq\omega\mu^{k}$ holds for all sufficiently large $k\in\mathbb{N}$, i.e., if the expression $\|x^{k}-x^{*}\|$ is dominated by a Q-linearly convergent null sequence.
-Let us fix a merely lower semicontinuous function $q\colon\mathbb{X}\to\overline{\mathbb{R}}$ and pick $x\in\operatorname{dom}q$ where $\operatorname{dom}q:=\{x\in\mathbb{X}\,|\,q(x)<\infty\}$ denotes the domain of $q$.
-Then the set
+Note that the R-linear convergence holds if there exist constants $a>0$ and $b\in(0,1)$ such that $\|x^{k}-x^{*}\|\leq ab^{k}$ holds for all sufficiently large $k\in\mathbb{N}$, in other words, $\|x^{k}-x^{*}\|$ is dominated by a Q-linearly convergent null sequence.
+Let a function $q\colon\mathbb{X}\to\overline{\mathbb{R}}$ be merely lower semicontinuous and let $x\in\operatorname{dom}q$ where $\operatorname{dom}q:=\{x\in\mathbb{X}\,|\,q(x)<\infty\}$ denotes the domain of $q$.
+Then
 ^$\widehat{\partial}q(x):=\left\{\eta\in\mathbb{X}\,\middle|\,\liminf\limits_{y\to x,\,y\neq x}\frac{q(y)-q(x)-\langle\eta,y-x\rangle}{\left\|y-x\right\|}\geq 0\right\}$
-is called the regular (or Fréchet) subdifferential of $q$ at $x$. Furthermore, the set
+is called the regular (or Fréchet) subdifferential of $q$ at $x$. Furthermore,
 $\partial q(x):=\left\{\eta\in\mathbb{X}\,\middle|\,\begin{aligned} &\exists\{x^{k}\},\{\eta^{k}\}\subset\mathbb{X}\colon\\ &\qquad x^{k}\to x,\,q(x^{k})\to q(x),\,\eta^{k}\to\eta,\,\eta^{k}\in\widehat{\partial}q(x^{k})\,\forall k\in\mathbb{N}\end{aligned}\right\}$
-is well known as the limiting (or Mordukhovich) subdifferential of $q$ at $x$.
-Clearly, we always have ^$\widehat{\partial}q(x)\subset\partial q(x)$ by construction of these sets.
+is denoted as the limiting (or Mordukhovich) subdifferential of $q$ at $x$.
+Naturally, ^$\widehat{\partial}q(x)\subset\partial q(x)$ always holds.
 Whenever $q$ is convex, both subdifferentials coincide, i.e.,
-^$\widehat{\partial}q(x)=\partial q(x)=\{\eta\in\mathbb{X}\,|\,\forall y\in\operatorname{dom}q\colon\,q(y)\geq q(x)+\langle\eta,y-x\rangle\}$
-is valid in this situation.
-By definition of the regular subdifferential, it is clear that whenever $x^{*}\in\operatorname{dom}q$ is a local minimizer of $q$, then $0\in\widehat{\partial}q(x^{*})$ hold by Fermat’s rule [22, Proposition 1.30(i)].
-Thus, the inclusion $0\in\partial q(x^{*})$ is a necessary optimality condition for $x^{*}$ being a local minimizer of $q$ as well.
+^$\widehat{\partial}q(x)=\partial q(x)=\{\eta\in\mathbb{X}\,|\,\forall y\in\operatorname{dom}q\colon\,q(y)\geq q(x)+\langle\eta,y-x\rangle\}.$
+By Fermat’s rule [23, Proposition 1.30(i)], whenever $x^{*}\in\operatorname{dom}q$ is a local minimizer of $q$, then $0\in\widehat{\partial}q(x^{*})$ hold, consequently $0\in\partial q(x^{*})$ holds as well.
 Since $f$ is continuously differentiabble, then for any fixed $x\in\operatorname{dom}\phi$, the sum rule
 $\partial(f+\phi)(x)=\nabla f(x)+\partial\phi(x)$ (2.1)
-holds, see [22, Proposition 1.30(ii)]. Fermat’s rule shows that the optimality condition
+holds, see [23, Proposition 1.30(ii)]. Fermat’s rule shows that the optimality condition
 $0\in\nabla f(x^{*})+\partial\phi(x^{*})$
-holds at any local minimizer $x^{*}\in\operatorname{dom}\phi$ of the composite optimization problem (Q).
-Any point $x^{*}\in\operatorname{dom}\phi$ satisfying this necessary optimality condition will be called an M-stationary point of (Q).
-We next introduce the famous Kurdyka–Łojasiewicz property which plays a central role in this manuscript.
-The version of this property stated below is a generalization of the classical Kurdyka–Łojasiewicz inequality for nonsmooth functions as introduced in [2, 5, 6] and afterwards used in the local convergence analysis of several nonsmooth optimization methods, see [1, 3, 7, 9, 10, 23, 24] for a couple of examples.
+holds at any local minimizer $x^{*}\in\operatorname{dom}\phi$ of (Q). Any point $x^{*}\in\operatorname{dom}\phi$ satisfying this necessary optimality condition is called an M-stationary point of (Q).
+We next introduce the Kurdyka–Łojasiewicz (KL) property which plays a vital role in this manuscript.
+The definition stated below is a generalization of the classical KL inequality for nonsmooth functions as introduced in [2, 5, 6] and afterwards used in the convergence analysis of many nonsmooth optimization algorithms, see [1, 3, 7, 10, 11, 24, 25] for a couple of examples.
 Definition 2.1.
 Let $g\colon\mathbb{X}\to\overline{\mathbb{R}}$ be lower semicontinuous.
-We say that $g$ has the KL property, where KL abbreviates Kurdyka–Łojasiewicz, at $x^{*}\in\{x\in\mathbb{X}\,|\,\partial g(x)\neq\emptyset\}$ if there exist a constant $\eta>0$, a neighborhood $U\subset\mathbb{X}$ of $x^{*}$, and a continuous concave function $\chi\colon[0,\eta]\to[0,\infty)$ which is continuously differentiable on $(0,\eta)$ and satisfies $\chi(0)=0$ as well as $\chi^{\prime}(t)>0$ for all $t\in(0,\eta)$ such that the so-called KL inequality
+We say that $g$ has the KL property at $x^{*}\in\{x\in\mathbb{X}\,|\,\partial g(x)\neq\emptyset\}$ if there exist a constant $\eta>0$, a neighborhood $U\subset\mathbb{X}$ of $x^{*}$, and a continuous concave function $\chi\colon[0,\eta]\to[0,\infty)$ which is continuously differentiable on $(0,\eta)$ and satisfies $\chi(0)=0$ as well as $\chi^{\prime}(t)>0$ for all $t\in(0,\eta)$ such that the so-called KL inequality
 $\chi^{\prime}\big{(}g(x)-g(x^{*})\big{)}\operatorname{dist}\big{(}0,\partial g(x)\big{)}\geq 1$
-holds for all $x\in U\cap\big{\{}x\in\mathbb{X}\,|\,g(x^{*})<g(x)<g(x^{*})+\eta\big{\}}$. The function $\chi$ from above is referred to as the desingularization function.
+holds for all $x\in U\cap\big{\{}x\in\mathbb{X}\,|\,g(x^{*})<g(x)<g(x^{*})+\eta\big{\}}$. The function $\chi$ is denoted as the desingularization function.
 A popular example of the desingularization function is given by $\chi(t):=ct^{\theta}$ for $\theta\in(0,1]$ and some constant $c>0$, where the parameter $\theta$ is called the KL exponent, see [6, 20].
-3 NPG with Average Line Search and Convergence Results
-We first propose the nonmonotone proximal gradient method with average line search in Algorithm 3.1.
-Algorithm 3.1 Nonmonotone Proximal Gradient Method with Average Line Search
+3 The NPG Method with Average Line Search and Convergence Results
+We first recall the nonmonotone proximal gradient method with average line search in Algorithm 3.1.
+Algorithm 3.1 A Nonmonotone Proximal Gradient Method with Average Line Search
 0:  $\tau>1$, $0<\gamma_{\min}\leq\gamma_{\max}<\infty$, $\delta\in(0,1)$, $\frac{4}{5}\leq p_{\min}\leq 1$, $x^{0}\in\operatorname{dom}q$.
@@
 9:  return  $x^{k}$
-In order to guarantee the convergence of Algorithm 3.1, we need the following requirement.
+In order to guarantee the convergence of Algorithm 3.1, we introduce the following requirements.
 Assumption 3.2.
@@
     The function $\nabla f\colon\mathbb{X}\to\mathbb{X}$ is locally Lipschitz continuous.
-Note that Assumption 3.2 (a) and (b) are the same as [18, Assumption 3.1].
-Assumption 3.2 (a) guarantees that the optimization problems (Q) are solvable.
-Assumption 3.2 (b) implies that $g$ is coercive, which consequently implies that subproblems (3.1) always have solutions, and hence the algorithm is well-defined.
-Combination of [18, Lemma 3.1] and [12, Lemma 4.1] implies that Algorithm 3.1 is well-defined. The following results are fundamental for the convergence analysis.
+Note that Assumption 3.2 (a) and (b) are the same as those presented in [19, Assumption 3.1].
+Assumption 3.2 (a) guarantees that the optimization problem (Q) is solvable.
+Assumption 3.2 (b) implies that $g$ is coercive, which consequently indicates that subproblems (3.1) always have solutions, and hence the algorithm is well-defined.
+The combination of [19, Lemma 3.1] and [13, Lemma 4.1] implies that Algorithm 3.1 is well-defined. The following results are fundamental to the convergence analysis.
 Proposition 3.3.
@@
   2. (b)
-    Every iterate $x^{k}$ remains in the sublevel set $\mathcal{L}_{q}(x^{0})\subset\operatorname{dom}g$.
+    Every iterate $x^{k}$ remains in the sublevel set $\mathcal{L}_{q}(x^{0}):=\{x\in\operatorname{dom}q\,|\,q(x)\leq q(x^{0})\}\subset\operatorname{dom}g$.
   3. (c)
@@
 Proof.
-Note that (a)$\sim$(d) have been observed in [12, Lemma 4.2] and [12, Lemma 4.3], so we omit the proof. It remains to consider (e).
+Note that (a)--(d) have been observed in [13, Lemma 4.2] and [13, Lemma 4.3], so we omit the proof. It remains to consider (e).
 When $k=0$, one has $\Phi_{0}=q(x^{0})$ by the initialization in Algorithm 3.1, then for any $k>0$, from (3.2), Steps 5 and 6, one has
@@
 We immediately obtain that $\|\hat{x}^{k}-x^{k}\|\to_{\mathcal{K}}0$ (otherwise, the left-hand side in (3.5) goes to infinity) and consequently $\hat{x}^{k}\to_{\mathcal{K}}\bar{x}$ from Assumption 3.2 (b).
-By Assumption 3.2 (a), using the mean-value theorem yields the existence of a point $\xi^{k}$ on the line segment connecting $x^{k}$ with $\hat{x}^{k}$ such that
+By Assumption 3.2 (a), the mean-value theorem yields that there exists $\xi^{k}$ on the line segment connecting $x^{k}$ with $\hat{x}^{k}$ such that
 $q(\hat{x}^{k})-q(x^{k})$ $=f(\hat{x}^{k})+g(\hat{x}^{k})-f(x^{k})-g(x^{k})$
 $=\langle\nabla f(\xi^{k}),\hat{x}^{k}-x^{k}\rangle+g(\hat{x}^{k})-g(x^{k}).$
-Substituting the resulting expression for $g(\hat{x}^{k})-g(x^{k})$ into (3.5) yields
+Recall (3.5), we have
 $\langle\nabla f(x^{k})-\nabla f(\xi^{k}),\hat{x}^{k}-x^{k}\rangle+\frac{\hat{\gamma}_{k}}{2}\|\hat{x}^{k}-x^{k}\|^{2}+q(\hat{x}^{k})-q(x^{k})\leq 0$ (3.6)
@@
 This completes the proof. ∎
-By means of the proof of Lemma 3.4 and the technique of [15, Lemma 4.1], we immediately have the following desired result, i.e., around the accumulation point $\bar{x}$ of $\{x^{k}\}_{k\in\mathbb{N}}$, the associated stepsize sequence remains (uniformly) bounded.
-This plays a central role in the convergence of entire sequence.
-We omit the proof because it is highly similar with [17, Lemma 4.1].
+By means of the proof of Lemma 3.4 and the technique outlined in [16, Lemma 4.1], we immediately obtain the following desired result, i.e., around the accumulation point $\bar{x}$ of $\{x^{k}\}_{k\in\mathbb{N}}$, the associated stepsize sequence remains (uniformly) bounded.
+We omit the proof because it closely resembles that of [18, Lemma 4.1].
 Lemma 3.5.
-Let Assumption 3.2 hold, $\{x^{k}\}_{k\in\mathbb{N}}$ be any sequence generated by Algorithm 3.1, and $\bar{x}$ is an accumulation point of $\{x^{k}\}_{k\in\mathbb{N}}$.
-Then, for any $\rho>0$, there is a constant $\bar{\gamma}_{\rho}>0$ (usually depending on $\rho$) such that $\gamma_{k}\leq\bar{\gamma}_{\rho}$ for all $k\in\mathbb{N}$ such that $x^{k}\in B_{\rho}(\bar{x})$.
-The following result provides the main global (subsequential) convergence of Algorithm 3.1.
+Let Assumption 3.2 hold, $\{x^{k}\}_{k\in\mathbb{N}}$ be any sequence generated by Algorithm 3.1, and $\bar{x}$ be its accumulation point.
+Then, for any $\rho>0$, there is a constant $\bar{\gamma}_{\rho}>0$ (depending on $\rho$) such that $\gamma_{k}\leq\bar{\gamma}_{\rho}$ for all $k\in\mathbb{N}$ such that $x^{k}\in B_{\rho}(\bar{x})$.
+The following result refers to the subsequential convergence of Algorithm 3.1.
 Theorem 3.6.
@@
 So, it remains to prove $g(x^{k+1})\to_{\mathcal{K}}g(\bar{x})$.
-From the fact that $g$ is lower semicontinuous, one then has
+From the fact that $g$ is lower semicontinuous, then one has
 $g(\bar{x})\leq\liminf_{k\to_{\mathcal{K}}\infty}g(x^{k+1})\leq\limsup_{k\to_{\mathcal{K}}\infty}g(x^{k+1}).$ (3.7)
@@
 ∎
-Theorem 3.6 actually obtained the subsequential convergence, we next aim to give the convergence of the entire sequence.
+Theorem 3.6 actually obtained the subsequential convergence, we next aim to give the convergence of the entire sequence in the presence of the KL property.
 Let sufficiently small $\eta>0$ be the corresponding constant from the definition of the associated desingularization function $\chi$.
@@
 $C_{\rho}:=B_{\rho}(\bar{x}).$ (3.11)
-Finally, throughout the section, let $L_{\rho}>0$ be a (global) Lipschitz constant of $\nabla f$ on $C_{\rho}$ from (3.11). In view of Lemma 3.5, one has
+Finally, let $L_{\rho}>0$ be the Lipschitz constant of $\nabla f$ on $C_{\rho}$ from (3.11). In view of Lemma 3.5, one has
 $\gamma_{k}\leq\bar{\gamma}_{\rho}\quad\forall x^{k}\in C_{\rho}$ (3.12)
@@
 $\alpha:=\|x^{k_{0}}-\bar{x}\|+\frac{1}{l}\sqrt{\frac{2}{\delta p_{\min}\gamma_{\min}}}\sqrt{\Phi_{k_{0}}-q(\bar{x})}+\frac{\bar{\gamma}_{\rho}+L_{\rho}}{2l}\chi\big{(}\Phi_{k_{0}}-q(\bar{x})\big{)}.$ (3.15)
-satisfies $\alpha<\frac{1}{2}$, where $\rho>0$, $\bar{\gamma}_{\rho}>0$, and $l>0$ are the constants defined in (3.10), (3.12), and (3.13), respectively, while $L_{\rho}>0$ is a Lipschitz constant of $\nabla f$ on $C_{\rho}$ from (3.11) and $\delta>0$ as well as $\gamma_{\min}>0$ are the parameters from Algorithm 3.1.
+satisfies $\alpha<\frac{1}{2}$, where $\rho>0$, $\bar{\gamma}_{\rho}>0$, and $l>0$ are the constants defined in (3.10), (3.12), and (3.13), respectively, while $L_{\rho}>0$ is the Lipschitz constant of $\nabla f$ on $C_{\rho}$ from (3.11) and $\delta>0$ as well as $\gamma_{\min}>0$ are the parameters from Algorithm 3.1.
 Proof.
 From $x^{k}\to_{\mathcal{K}}\bar{x}$, then there exists a sufficiently large $k_{0}\in\mathcal{K}$ satisfying $\Phi_{k_{0}}-q(\bar{x})\geq 0$ sufficiently small deduced by Proposition 3.3 and Theorem 3.6.
-Recall that that $q$ satisfies KL property at $\bar{x}$, the corresponding disingularizatioon function is denoted as $\chi$.
-For such $\chi$, one has $\chi(0)=0$ and $\chi$ is continuous and increasing on its domain, therefore one knows $\chi\big{(}\Phi_{k_{0}}-q(\bar{x})\big{)}\geq 0$ sufficiently small.
+Recall that that $q$ satisfies the KL property at $\bar{x}$, the corresponding disingularizatioon function is denoted as $\chi$.
+For such $\chi$, one has $\chi(0)=0$ and $\chi$ is continuous and increasing on its domain, therefore $\chi\big{(}\Phi_{k_{0}}-q(\bar{x})\big{)}\geq 0$ is sufficiently small.
 Hence, each summand on the right-hand side (3.15) by taking an index $k_{0}\in\mathcal{K}$ sufficiently large yields $\alpha<\frac{1}{2}$.
 ∎
-Using these notations, we get the following result. We omit the proof, since it is highly similar with [17, Lemma 4.3].
+Using these notations, we obtain the following result. The proof is omitted because it is similar to [18, Lemma 4.3].
 Lemma 3.8.
@@
 holds for all sufficiently large $k\geq\hat{k}$ such that $x^{k}\in B_{\alpha}(\bar{x})$, where $\alpha$ denotes the constant from (3.15), $\bar{\gamma}_{\rho}>0$ is the constant from (3.12), and $L_{\rho}>0$ is the Lipschitz constant of $\nabla f$ on $C_{\rho}$ from (3.11).
-By employing the above results, the following theorem demonstrates that the whole sequence $\{x^{k}\}_{k\in\mathbb{N}}$ generated by Algorithm 3.1 is convergent to its accumulation point $\bar{x}$, provided that $q$ satisfies the KL property at $\bar{x}$.
+By employing the above results, the following theorem demonstrates that the whole sequence $\{x^{k}\}_{k\in\mathbb{N}}$ generated by Algorithm 3.1 convergences to its accumulation point $\bar{x}$, provided that $q$ satisfies the KL property at $\bar{x}$.
 Theorem 3.9.
@@
 By Proposition 3.3 (a) and (c), one knows that the whole sequence $\{\Phi_{k}\}_{k\in\mathbb{N}}$ is monotonically decreasing and converges to $q(\bar{x})$ from Theorem 3.6.
-This implies that $\Phi_{k}\geq q(\bar{x})$ holds for all $k\in\mathbb{N}$.
+It implies that $\Phi_{k}\geq q(\bar{x})$ holds for all $k\in\mathbb{N}$.
 If $\Phi_{k}=q(\bar{x})$ holds for some index $k\in\mathbb{N}$, which, by monotonicity, implies that $\Phi_{k+1}=q(\bar{x})$.
@@
 which says that $x^{k+1}=x^{k}$ holds.
-Since, by assumption, the subsequence $\{x^{k}\}_{\mathcal{K}}$ converges to $\bar{x}$, this implies that $x^{k}=\bar{x}$ for all $k\in\mathbb{N}$ sufficiently large.
-In particular, one has convergence of the entire (eventually constant) sequence $\{x^{k}\}_{k\in\mathbb{N}}$ to $\bar{x}$ in this situation.
+Since the subsequence $\{x^{k}\}_{\mathcal{K}}$ are assumed to converge to $\bar{x}$, this implies that $x^{k}=\bar{x}$ for all $k\in\mathbb{N}$ sufficiently large.
+Consequently, the entire (eventually constant) sequence $\{x^{k}\}_{k\in\mathbb{N}}$ converges to $\bar{x}$ in this situation.
 It remains to consider the case where $\Phi_{k}>q(\bar{x})$ for all $k\in\mathbb{N}$.
-Let $\alpha$ be the constant from (3.15), and $k_{0}\in\mathcal{K}$ be the corresponding iteration index which is used in the definition of $\alpha$, see Lemma 3.7.
+Let $\alpha$ be the constant from (3.15), and $k_{0}\in\mathcal{K}$ be the corresponding iteration index used in the definition of $\alpha$, see Lemma 3.7.
 One then has $0<\Phi_{k}-q(\bar{x})\leq\Phi_{k_{0}}-q(\bar{x})$ for all $k\geq k_{0}$.
@@
 $\Phi_{k_{0}}<q(\bar{x})+\eta.$ (3.17)
-Let $\chi\colon[0,\eta]\to[0,\infty)$ be the desingularization function which comes along with the validity of the KL property of $q$ at $\bar{x}$.
+Let $\chi\colon[0,\eta]\to[0,\infty)$ be the desingularization function which validates the KL property of $q$ at $\bar{x}$.
 Due to $\chi(0)=0$ and $\chi^{\prime}(t)>0$ for all $t\in(0,\eta)$, one obtains
@@
 Suppose that both statements hold for some $k\geq k_{0}$.
-Using the triangle inequality, the induction hypothesis, and the definition of $\alpha$, one obtains
+By the triangle inequality, the induction hypothesis, and the definition of $\alpha$, one obtains
 $\|x^{k+1}-\bar{x}\|\leq\sum_{i=k_{0}}^{k}\|x^{i+1}-x^{i}\|+\|x^{k_{0}}-\bar{x}\|\leq\alpha,$
@@
 $q(\bar{x})<\Phi_{i+1}<q(x^{i})\leq\Phi_{i}\leq\Phi_{k_{0}}<q(\bar{x})+\eta\quad\forall i\geq k_{0}\ \text{and}\ i\in\overline{S}.$ (3.22)
-Since $q$ has the KL property at $\bar{x}$, one has
+Recall that $q$ has the KL property at $\bar{x}$, one has
 $\chi^{\prime}\big{(}q(x^{i})-q(\bar{x})\big{)}\operatorname{dist}\big{(}0,\partial q(x^{i})\big{)}\geq 1\quad\forall i\geq k_{0}\ \text{and}\ i\in\overline{S}.$ (3.23)
@@
 Hence, statement (b) holds for $k+1$ in place of $k$, and this completes the induction.
-In particular, it follows from (a) that $x^{k}\in B_{\alpha}(\bar{x})$ for all $k\geq k_{0}$.
-And taking $k\to\infty$ in (3.19) therefore shows that $\{x^{k}\}_{k\in\mathbb{N}}$ is a Cauchy sequence and, thus, convergent.
-Since one already knows that $\bar{x}$ is an accumulation point, it follows that the entire sequence $\{x^{k}\}_{k\in\mathbb{N}}$ converges to $\bar{x}$.
+It follows from (a) that $x^{k}\in B_{\alpha}(\bar{x})$ for all $k\geq k_{0}$.
+Taking $k\to\infty$ in (3.19) shows that $\{x^{k}\}_{k\in\mathbb{N}}$ is a Cauchy sequence and hence convergent.
+Since $\bar{x}$ is an accumulation point, then the entire sequence $\{x^{k}\}_{k\in\mathbb{N}}$ converges to $\bar{x}$.
 ∎
-Note that if all index $k\geq k_{0}$ belong to $S$ or if $p_{k}:=1$, then Algorithm 3.1 degenerates into monotone proximal gradient, see [18, Algorithm 3.1] or [15, Algorithm 3.1], and the corresponding rate-of-convergence result was obtained in [14].
-So, we in the following state the rate-of-convergence result of the nonmonotone proximal gradient methods (in general case), readers may find more details for the proof in [17, Theorem 4.6].
+Note that if all indices $k\geq k_{0}$ belong to $S$ or if $p_{k}:=1$, then Algorithm 3.1 degenerates into a monotone proximal gradient method, see [19, Algorithm 3.1] or [16, Algorithm 3.1], and the corresponding rate-of-convergence result was obtained in [15].
+So, we in the following state the rate-of-convergence result of the NPG method (in the general case).
+Readers may find more details for the proof in [18, Theorem 4.6].
 Theorem 3.10.
@@
     for sufficiently large $k\geq k_{0}$.
-4 NPG with Max Line Search and Convergence Analysis
-This section focuses on another type of nonmonotone proximal gradient method, where the line search is chosen as max-type rule.
-Let us now introduce the algorithm which is exactly from [18, Algorithm 4.1].
-Algorithm 4.1 Nonmonotone Proximal Gradient Method with Max Line Search
+4 The NPG method with Max Line Search and Convergence Analysis
+This section focuses on another type of the nonmonotone proximal gradient method, where the line search is chosen as the max-type rule.
+Let us now introduce the algorithm which is derived from [19, Algorithm 4.1].
+Algorithm 4.1 A Nonmonotone Proximal Gradient Method with Max Line Search
 0:  $\tau>1$, $0<\gamma_{\min}\leq\gamma_{\max}<\infty$, $m\in\mathbb{N}$, $\delta\in(0,1)$, $x^{0}\in\operatorname{dom}q$.
@@
 8:  return  $x^{k}$
-Throughout the section, let us denote $l(k)\in\{k-m_{k},\ldots,k\}$ an index satisfying
+Throughout this section, let us define $l(k)\in\{k-m_{k},\ldots,k\}$ as an index satisfying
 $q(x^{l(k)})=\max_{j=0,1,\ldots,m_{k}}q(x^{k-j})$
@@
 Note that Assumption 3.2 (b) and Assumption 4.2 imply that $q$ is uniformly continuous on the corresponding sublevel $\mathcal{L}_{q}(x^{0})$.
-The requirement of uniformly continuity about objective function plays an essential role in the context of nonmonotone line search rules with maximum taste [13].
-We here recall the results again for the convenience of readers.
+The requirement of uniform continuity about the objective function plays an essential role in the context of nonmonotone line search rules with the maximum taste [14].
+We here recall some results for the convenience of readers.
 Proposition 4.3.
-Let Assumption 3.2 and Assumption 4.2 hold and let $\{x^{k}\}$ be any sequence generated by Algorithm 4.1, then
+Let Assumption 3.2 and Assumption 4.2 hold and let $\{x^{k}\}_{k\in\mathbb{N}}$ be any sequence generated by Algorithm 4.1, then
   1. (a)
-    the sequence $\{q(x^{l(k)})\}$ is monotonically decreasing,
+    the sequence $\{q(x^{l(k)})\}_{k\in\mathbb{N}}$ is monotonically decreasing,
   2. (b)
-    the sequences $\{x^{k}\}$, $\{x^{l(k)}\}$ $\subset\mathcal{L}_{q}(x^{0})\subset\operatorname{dom}g$,
+    the sequences $\{x^{k}\}_{k\in\mathbb{N}}$, $\{x^{l(k)}\}_{k\in\mathbb{N}}$ $\subset\mathcal{L}_{q}(x^{0})\subset\operatorname{dom}g$,
   3. (c)
@@
   4. (d)
-    suppose $\bar{x}$ is an accumulation point of $\{x^{k}\}$ such that $x^{k}\to_{\mathcal{K}}\bar{x}$ holds.
+    suppose $\bar{x}$ is an accumulation point of $\{x^{k}\}_{k\in\mathbb{N}}$ such that $x^{k}\to_{\mathcal{K}}\bar{x}$ holds.
 Then $\bar{x}$ is an M-stationary point of (Q), and $\gamma_{k}\|x^{k+1}-x^{k}\|\to_{\mathcal{K}}0$ is valid.
-With the aid of Proposition 4.3 (d) as well as the techniques in Lemma 3.5, we can obtain the following results.
+With the aid of Proposition 4.3 (d) as well as the techniques in Lemma 3.5, we can obtain the following result.
 Lemma 4.4.
-Let Assumption 3.2 and Assumption 4.2 hold, let $\{x^{k}\}$ be any sequence generated by Algorithm 4.1, $\bar{x}$ is an accumulation point of $\{x^{k}\}$ and consequently let $\{x^{k}\}_{\mathcal{K}}$ be a subsequence converging to some point $\bar{x}$.
-Then, for any $\rho>0$, there is a constant $\bar{\gamma}_{\rho}>0$ (usually depending on $\rho$) such that $\gamma_{k}\leq\bar{\gamma}_{\rho}$ for all $k\in\mathbb{N}$ such that $x^{k}\in B_{\rho}(\bar{x})$.
-We now discuss the convergence properties of the sequence generated by Algorithm 4.1 in the presence of KL property of $q$ at the mentioned accumulation point $\bar{x}$.
+Let Assumption 3.2 and Assumption 4.2 hold, let $\{x^{k}\}_{k\in\mathbb{N}}$ be any sequence generated by Algorithm 4.1, $\bar{x}$ is an accumulation point of $\{x^{k}\}_{k\in\mathbb{N}}$ and consequently let $\{x^{k}\}_{k\in\mathcal{K}}$ be a subsequence converging to some point $\bar{x}$.
+Then, for any $\rho>0$, there is a constant $\bar{\gamma}_{\rho}>0$ (depending on $\rho$) such that $\gamma_{k}\leq\bar{\gamma}_{\rho}$ for all $k\in\mathbb{N}$ such that $x^{k}\in B_{\rho}(\bar{x})$.
+We now discuss the convergence properties of the sequence generated by Algorithm 4.1 in the presence of the KL property of $q$ at the mentioned accumulation point $\bar{x}$.
 Let sufficiently small $\eta>0$ be the corresponding constant from the definition of the associated desingularization function $\chi$.
@@
 $C_{\rho}:=B_{2\rho}(\bar{x})\cap\mathcal{L}_{q}(x^{0}).$ (4.6)
-Consequently, let $L_{\rho}>0$ be a (global) Lipschitz constant of $\nabla f$ on $C_{\rho}$ from (4.6). In view of Lemma 4.4, let $\bar{\gamma}_{\rho}>0$ be satisfying
+Consequently, let $L_{\rho}>0$ be the (global) Lipschitz constant of $\nabla f$ on $C_{\rho}$ from (4.6). In view of Lemma 4.4, let $\bar{\gamma}_{\rho}>0$ satisfy
 $\gamma_{k}\leq\bar{\gamma}_{\rho}\quad\forall x^{k}\in C_{\rho}.$ (4.7)
-Thanks to the above notation, we can find a new neighborhood centered at the accumulation point, whose radium is very small which is mainly used to to ensure the convergence of the whole sequence under the assumption of KL property of $q$, we now introduce such radium.
+Proposition 4.5.
+Let Assumption 3.2 and Assumption 4.2 hold, and let $\{x^{k}\}_{k\in\mathbb{N}}$ be any sequence generated by Algorithm 4.1. Then there exists a constant $c_{\hat{k}}>0$ such that
+$\|x^{l(k+1)}-x^{k+1}\|\leq c_{\hat{k}}\|x^{l(k+1)}-x^{l(k+1)-1}\|\quad\forall k\geq\hat{k},$
+where $\hat{k}$ is defined in (4.3).
+Proof.
+By contradiction, for arbitrary $c>0$, there exists a $\bar{k}\geq\hat{k}$ such that
+$\|x^{l(\bar{k}+1)}-x^{\bar{k}+1}\|>c\|x^{l(\bar{k}+1)}-x^{l(\bar{k}+1)-1}\|.$
+Meanwhile, for the sufficiently large $\bar{k}\geq\hat{k}$, from (4.3), one has
+$\|x^{l(\bar{k}+1)}-x^{\bar{k}+1}\|\leq\sum_{i=\bar{k}+1-m}^{\bar{k}+1}\|x^{i+1}-x^{i}\|\leq m\eta.$
+Therefore, one has
+$m\eta>c\|x^{l(\bar{k}+1)}-x^{l(\bar{k}+1)-1}\|.$
+Now, choose $c:=\frac{1}{\|x^{l(\bar{k}+1)}-x^{l(\bar{k}+1)-1}\|}$, we obtain:
+$m\eta>1,$
+which contradicts the fact that $\eta$ is sufficiently small and $m>0$ is a fixed number. ∎
+Thanks to the above notation, we can identify a new neighborhood centered at the accumulation point, with a very small radius, which is primarily used to to ensure the convergence of the whole sequence under the assumption of the KL property of $q$, we now introduce such radius.
 Lemma 4.6.
-Let Assumption 3.2 and Assumption 4.2 hold, and let $\{x^{k}\}$ be any sequence generated by Algorithm 4.1.
-Suppose that $\{x^{k}\}_{\mathcal{K}}$ is a subsequence converging to some limit point $\bar{x}$, and that $q$ has the KL property at $\bar{x}$ with desingularization function $\chi$.
+Let Assumption 3.2 and Assumption 4.2 hold, and let $\{x^{k}\}_{k\in\mathbb{N}}$ be any sequence generated by Algorithm 4.1.
+Suppose that $\{x^{k}\}_{k\in\mathcal{K}}$ is a subsequence converging to some limit point $\bar{x}$, and that $q$ has the KL property at $\bar{x}$ with desingularization function $\chi$.
 Then there is a sufficiently large constant $k_{0}\in\mathcal{K}\cap\overline{K}$ such that $k_{0}-m-1$ is also sufficiently large, where $K$ is defined by (4.4), then the corresponding constant is defined as
 $\alpha$ $:=\|x^{k_{0}}-\bar{x}\|$ (4.8)
-$\leq\left(\frac{8m(\bar{\gamma}_{\rho}+L_{\rho})}{3\delta\gamma_{\min}}\sqrt{\frac{\bar{\gamma}_{\rho}+L_{\rho}}{\mu}}+\frac{m}{3}\sqrt{\frac{2(\bar{\gamma}_{\rho}+L_{\rho})}{\delta\gamma_{\min}-\mu}}\left(\frac{3}{m}+\frac{2(\bar{\gamma}_{\rho}+L_{\rho})}{\delta\gamma_{\min}}\right)\right)\chi\big{(}q(x^{l(k_{0})})-q(\bar{x})\big{)}$
-$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ +\left(\frac{4m}{3}\sqrt{\frac{\bar{\gamma}_{\rho}+L_{\rho}}{\mu}}+\frac{m}{3}\sqrt{\frac{2(\bar{\gamma}_{\rho}+L_{\rho})}{\delta\gamma_{\min}-\mu}}+\sqrt{\frac{\delta\gamma_{\min}}{\delta\gamma_{\min}-\mu}}\right)\sqrt{\frac{2\big{(}q(x^{l(k_{0}-m-1)})-q(\bar{x})\big{)}}{\delta\gamma_{\min}}}$
+$\leq\left(\frac{8\sqrt{2}m\left(c_{\hat{k}}+1\right)(\bar{\gamma}_{\rho}+L_{\rho})}{3\delta\gamma_{\min}}\sqrt{\frac{\bar{\gamma}_{\rho}+L_{\rho}}{\mu}}+\frac{m}{3}\sqrt{\frac{2(\bar{\gamma}_{\rho}+L_{\rho})}{\delta\gamma_{\min}-\mu}}\left(\frac{3}{m}+\frac{2(\bar{\gamma}_{\rho}+L_{\rho})}{\delta\gamma_{\min}}\right)\right)$
+$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \cdot\chi\big{(}q(x^{l(k_{0})})-q(\bar{x})\big{)}$
+$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ +\left(\frac{4\sqrt{2}m\left(c_{\hat{k}}+1\right)}{3}\sqrt{\frac{\bar{\gamma}_{\rho}+L_{\rho}}{\mu}}+\frac{m}{3}\sqrt{\frac{2(\bar{\gamma}_{\rho}+L_{\rho})}{\delta\gamma_{\min}-\mu}}+\sqrt{\frac{\delta\gamma_{\min}}{\delta\gamma_{\min}-\mu}}\right)$
+$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \cdot\sqrt{\frac{2\big{(}q(x^{l(k_{0}-m-1)})-q(\bar{x})\big{)}}{\delta\gamma_{\min}}}$
 satisfies $\alpha<\frac{1}{2}$, where $\rho>0$, $\bar{\gamma}_{\rho}>0$, $\mu>0$ are the constants defined in (4.5), (4.7), and (4.4), respectively, while $\delta>0$ and $\gamma_{\min}>0$ are the parameters from Algorithm 4.1.
 Proof.
-From the fact that $\{x^{k}\}_{\mathcal{K}}\to\bar{x}$, the decreasement of $\{q(x^{l(k)})\}$, we have $\|x^{k_{0}}-\bar{x}\|$ is sufficiently small by the sufficiently large $k_{0}\in\mathcal{K}$, and $q(x^{l(k)})-q(\bar{x})\to 0$ from Assumption 4.2.
+From the fact that $x^{k}\to_{\mathcal{K}}\bar{x}$, the decrease of $\{q(x^{l(k)})\}_{k\in\mathbb{N}}$, we have $\|x^{k_{0}}-\bar{x}\|$ is sufficiently small by the sufficiently large $k_{0}\in\mathcal{K}$, and $q(x^{l(k)})-q(\bar{x})\to 0$ from Assumption 4.2.
 Then we have $\chi(q(x^{l(k)})-q(\bar{x}))\to 0$ from the continuity of $\chi$ and $\chi(0)=0$.
@@
 ∎
-Taking the similar analysis with Lemma 3.8, we obtain the following result.
+Taking the similar analysis to Lemma 3.8, we obtain the following result.
 Lemma 4.7.
-Let Assumption 3.2 and Assumption 4.2 hold, and let $\{x^{k}\}$ be any sequence generated by Algorithm 4.1.
-Suppose that $\{x^{k}\}_{\mathcal{K}}$ is a subsequence converging to some limit point $\bar{x}$, then
+Let Assumption 3.2 and Assumption 4.2 hold, and let $\{x^{k}\}_{k\in\mathbb{N}}$ be any sequence generated by Algorithm 4.1.
+Suppose that $\{x^{k}\}_{k\in\mathcal{K}}$ is a subsequence converging to some limit point $\bar{x}$, then
 $\operatorname{dist}\big{(}0,\partial q(x^{k+1})\big{)}\leq\big{(}\bar{\gamma}_{\rho}+L_{\rho}\big{)}\|x^{k+1}-x^{k}\|$
@@
 Theorem 4.8.
-Let Assumption 3.2 and Assumption 4.2 hold, and let $\{x^{k}\}$ be any sequence generated by Algorithm 4.1.
-Suppose that $\{x^{k}\}_{\mathcal{K}}$ is a subsequence converging to some limit point $\bar{x}$, and that $q$ has the KL property at $\bar{x}$.
+Let Assumption 3.2 and Assumption 4.2 hold, and let $\{x^{k}\}_{k\in\mathbb{N}}$ be any sequence generated by Algorithm 4.1.
+Suppose that $\{x^{k}\}_{k\in\mathcal{K}}$ is a subsequence converging to some limit point $\bar{x}$, and that $q$ has the KL property at $\bar{x}$.
 If $q(x^{l(k)})\neq q(\bar{x})$ for all $k\in\mathbb{N}$ and for any $v\geq k_{0}$ where $k_{0}$ is used to define $\alpha$ in Lemma 4.6, if $x^{k}\in B_{\alpha}(\bar{x})$ holds for all $k_{0}\leq k\leq v$, then we have
@@
 $q(\bar{x})<q(x^{l(k)})\leq q(x^{l(k_{0})})<q(\bar{x})+\eta\quad\forall k\geq k_{0}.$ (4.9)
-Let $\chi:[0,\eta]\to[0,\infty)$ be the desingularization function which comes along with the validity of the KL property of $q$.
-Due to $\chi(0)=0$ and $\chi^{\prime}(t)>0$ for all $t\in(0,\eta)$, one obtains
+Let $\chi:[0,\eta]\to[0,\infty)$ be the desingularization function which validates of the KL property of $q$. Due to $\chi(0)=0$ and $\chi^{\prime}(t)>0$ for all $t\in(0,\eta)$, one obtains
 $\chi\big{(}q(x^{l(k)})-q(\bar{x})\big{)}\geq 0\quad\forall k\geq k_{0}.$
@@
 Theorem 4.9.
-Let Assumption 3.2 and Assumption 4.2,hold, and let $\{x^{k}\}$ be any sequence generated by Algorithm 4.1.
-Suppose that $\{x^{k}\}_{\mathcal{K}}$ is a subsequence converging to some limit point $\bar{x}$, and that $q$ has the KL property at $\bar{x}$, then the entire sequence $\{x^{k}\}$ converges to $\bar{x}$.
+Let Assumption 3.2 and Assumption 4.2 hold, and let $\{x^{k}\}_{k\in\mathbb{N}}$ be any sequence generated by Algorithm 4.1.
+Suppose that $\{x^{k}\}_{k\in\mathcal{K}}$ is a subsequence converging to some limit point $\bar{x}$, and that $q$ has the KL property at $\bar{x}$, then the entire sequence $\{x^{k}\}_{k\in\mathbb{N}}$ converges to $\bar{x}$.
 Proof.
 By Proposition 4.3 (a), one knows that the whole sequence $\{q(x^{l(k)})\}$ is monotonically decreasing and convergent to $q(\bar{x})$ by assumption.
-This implies that $q(x^{l(k)})\geq q(\bar{x})$ for all $k\in\mathbb{N}$.
-If $q(x^{l(\bar{k})})=q(\bar{x})$ holds for some index $\bar{k}\in\mathbb{N}$, which, by monotonicity, implies that $q(x^{l(\bar{k}+1)})=q(\bar{x})$, then we claim that $x^{k+1}=x^{k}$ for all $k\geq\bar{k}+m_{\bar{k}}$ ($m_{\bar{k}}=m$ if $\bar{k}$ is sufficiently large) by [25, Lemma 3.2 (iv)].
-Since, by assumption, the subsequence $\{x^{k}\}_{\mathcal{K}}$ converges to $\bar{x}$, this implies that $x^{k}=\bar{x}$ for all $k\in\mathbb{N}$ sufficiently large.
-In particular, we have the convergence of the entire (eventually constant) sequence $\{x^{k}\}$ to $\bar{x}$ in this situation.
+It implies that $q(x^{l(k)})\geq q(\bar{x})$ for all $k\in\mathbb{N}$.
+If $q(x^{l(\bar{k})})=q(\bar{x})$ holds for some index $\bar{k}\in\mathbb{N}$, which, by monotonicity, implies that $q(x^{l(\bar{k}+1)})=q(\bar{x})$, then we claim that $x^{k+1}=x^{k}$ for all $k\geq\bar{k}+m_{\bar{k}}$ ($m_{\bar{k}}=m$ if $\bar{k}$ is sufficiently large) by [26, Lemma 3.2 (iv)].
+Since the subsequence $\{x^{k}\}_{k\in\mathcal{K}}$ is assumed to converge to $\bar{x}$, this implies that $x^{k}=\bar{x}$ for all $k\in\mathbb{N}$ sufficiently large.
+Then the entire (eventually constant) sequence $\{x^{k}\}_{k\in\mathbb{N}}$ is convergent to $\bar{x}$ in this situation.
 It remains to consider the case where $q(x^{l(k)})>q(\bar{x})$ for all $k\in\mathbb{N}$.
@@
     $\sum_{i=k_{0}}^{k}\|x^{i+1}-x^{i}\|$ (4.14)
-    $\leq\left(\frac{8m(\bar{\gamma}_{\rho}+L_{\rho})}{3\delta\gamma_{\min}}\sqrt{\frac{\bar{\gamma}_{\rho}+L_{\rho}}{\mu}}+\frac{m}{3}\sqrt{\frac{2(\bar{\gamma}_{\rho}+L_{\rho})}{\delta\gamma_{\min}-\mu}}\left(\frac{3}{m}+\frac{2(\bar{\gamma}_{\rho}+L_{\rho})}{\delta\gamma_{\min}}\right)\right)\chi\big{(}q(x^{l(k_{0})})-q(\bar{x})\big{)}$
-    $\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ +\left(\frac{4m}{3}\sqrt{\frac{\bar{\gamma}_{\rho}+L_{\rho}}{\mu}}+\frac{m}{3}\sqrt{\frac{2(\bar{\gamma}_{\rho}+L_{\rho})}{\delta\gamma_{\min}-\mu}}+\sqrt{\frac{\delta\gamma_{\min}}{\delta\gamma_{\min}-\mu}}\right)\sqrt{\frac{2\big{(}q(x^{l(k_{0}-m-1)})-q(\bar{x})\big{)}}{\delta\gamma_{\min}}},$
+    $\leq\left(\frac{8\sqrt{2}m\left(c_{\hat{k}}+1\right)(\bar{\gamma}_{\rho}+L_{\rho})}{3\delta\gamma_{\min}}\sqrt{\frac{\bar{\gamma}_{\rho}+L_{\rho}}{\mu}}+\frac{m}{3}\sqrt{\frac{2(\bar{\gamma}_{\rho}+L_{\rho})}{\delta\gamma_{\min}-\mu}}\left(\frac{3}{m}+\frac{2(\bar{\gamma}_{\rho}+L_{\rho})}{\delta\gamma_{\min}}\right)\right)$
+    $\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \cdot\chi\big{(}q(x^{l(k_{0})})-q(\bar{x})\big{)}$
+    $\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ +\left(\frac{4\sqrt{2}m\left(c_{\hat{k}}+1\right)}{3}\sqrt{\frac{\bar{\gamma}_{\rho}+L_{\rho}}{\mu}}+\frac{m}{3}\sqrt{\frac{2(\bar{\gamma}_{\rho}+L_{\rho})}{\delta\gamma_{\min}-\mu}}+\sqrt{\frac{\delta\gamma_{\min}}{\delta\gamma_{\min}-\mu}}\right)$
+    $\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \cdot\sqrt{\frac{2\big{(}q(x^{l(k_{0}-m-1)})-q(\bar{x})\big{)}}{\delta\gamma_{\min}}}$
 where $\alpha$ is defined in (4.8).
@@
 Case 2: $k\in K$. For sufficiently large $k_{0}$, from (4.3), one has
-$\|x^{l(k+1)}-x^{k+1}\|\leq\sum_{i=k+1-m}^{k+1}\|x^{i+1}-x^{i}\|\leq m\eta,\quad\forall k\geq k_{0},$ (4.19)
-which means that $\|x^{l(k+1)}-x^{k+1}\|\to 0$. From (4.1), one has
-$\gamma_{k}(x^{k}-x^{k+1})+\nabla f(x^{k+1})-\nabla f(x^{k})\in\hat{\partial}q(x^{k+1})\quad\forall k\in\mathbb{N}.$
-Meanwhile, for all $k\in K$ where $K$ is defined in (4.4), we know that $x^{l(k+1)}\neq x^{k+1}$.
-We obtain that $\|x^{l(k+1)}-x^{k+1}\|\to 0$ by Theorem 4.7.
-Hence, by the definition of (Fréchet) subdifferential, we have
-$q(x^{k+1})-q(x^{l(k+1)})-$ (4.20)
-$\gamma_{l(k+1)-1}(x^{l(k+1)-1}-x^{l(k+1)})+\nabla f(x^{l(k+1)})-\nabla f(x^{l(k+1)-1}),x^{k+1}-x^{l(k+1)}\right>$
-$\geq o(\|x^{k+1}-x^{l(k+1)}\|)$
-Recall again that $x^{i}\in B_{\alpha}(\bar{x})$ for all $k_{0}\leq i\leq k+1$, then from (4.19), we have
+$\|x^{l(i+1)}-x^{i+1}\|\leq\sum_{j=i+1-m}^{i+1}\|x^{j+1}-x^{j}\|\leq m\eta,\quad\forall i\geq k_{0},$ (4.19)
+which means that $\|x^{l(i+1)}-x^{i+1}\|\to 0$. By (4.1), for each $i\in\mathbb{N}$, one has
+$\nabla f(x^{l(i+1)-1}),x^{l(i+1)}-x^{l(i+1)-1}\right>+\frac{\gamma_{l(i+1)-1}}{2}\|x^{l(i+1)}-x^{l(i+1)-1}\|^{2}+g(x^{l(i+1)})$
+$\nabla f(x^{l(i+1)-1}),x^{i+1}-x^{l(i+1)-1}\right>+\frac{\gamma_{l(i+1)-1}}{2}\|x^{i+1}-x^{l(i+1)-1}\|^{2}+g(x^{i+1}),$
+which implies that
+$q(x^{l(i+1)})-q(x^{i+1})$ $\nabla f(x^{l(i+1)-1}),x^{i+1}-x^{l(i+1)}\right>+f(x^{l(i+1)})-f(x^{i+1})$ (4.20)
+$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ +\frac{\gamma_{l(i+1)-1}}{2}\left(\|x^{i+1}-x^{l(i+1)-1}\|^{2}-\|x^{l(i+1)}-x^{l(i+1)-1}\|^{2}\right).$
+Recall again that $x^{i}\in B_{\alpha}(\bar{x})$ for all $k_{0}\leq i\leq k+1$, then $x^{i+1}\in C_{\rho}$ holds and from (4.19), we have
 $\|x^{l(i+1)}-\bar{x}\|$ $\leq\|x^{l(i+1)}-x^{i+1}\|+\|x^{i+1}-x^{i}\|+\|x^{i}-\bar{x}\|$
@@
 for all $k_{0}\leq i\leq k+1$.
-Therefore, for all $k_{0}\leq i\leq k+1$, we have $\gamma_{l(i+1)-1}\leq\bar{\gamma}_{\rho}$ from Lemma 4.4 that $\gamma_{l(i+1)-1}\leq\bar{\gamma}_{\rho}$, and $x^{l(i+1)-1},x^{l(i+1)}\in C_{\rho}$, then we have
-$\|\nabla f(x^{l(i+1)})-\nabla f(x^{l(i+1)-1})\|\leq L_{\rho}\|x^{l(i+1)}-x^{l(i+1)-1}\|.$
-Therefore, (4.20) implies that
-$q(x^{i+1})-q(x^{l(i+1)})+(\bar{\gamma}_{\rho}+L_{\rho})\|x^{l(i+1)-1}-x^{l(i+1)}\|\|x^{i+1}-x^{l(i+1)}\|\geq o(\|x^{i+1}-x^{l(i+1)}\|)$
-for all $k_{0}\leq i\leq k+1$ and $i\in K$. Employing the inequality $a^{2}+b^{2}\geq 2ab$ for all $a,b\in\mathbb{R}$ as well as $\|x^{i+1}-x^{l(i+1)}\|\to 0$ yields that
-$q(x^{i+1})-q(x^{l(i+1)})+\frac{\bar{\gamma}_{\rho}+L_{\rho}}{2}\left(\|x^{l(i+1)-1}-x^{l(i+1)}\|^{2}+\|x^{i+1}-x^{l(i+1)}\|^{2}\right)$
-$\geq\frac{\bar{\gamma}_{\rho}+L_{\rho}}{2}\|x^{i+1}-x^{l(i+1)}\|^{2}\quad\forall k_{0}\leq i\leq k+1\ \text{and}\ i\in K,$
-which implies that
-$\frac{\bar{\gamma}_{\rho}+L_{\rho}}{2}\|x^{l(i+1)-1}-x^{l(i+1)}\|^{2}\geq q(x^{l(i+1)})-q(x^{i+1})$ (4.21)
-for all $k_{0}\leq i\leq k+1$ and $i\in K$. By (4.21) and (4.4), we have
-$\frac{\mu}{2}\|x^{i+1}-x^{i}\|^{2}<q(x^{l(i+1)})-q(x^{i+1})\leq\frac{\bar{\gamma}_{\rho}+L_{\rho}}{2}\|x^{l(i+1)-1}-x^{l(i+1)}\|^{2}$
+Therefore, for all $k_{0}\leq i\leq k+1$, we have $\gamma_{l(i+1)-1}\leq\bar{\gamma}_{\rho}$ from Lemma 4.4, and $x^{i},x^{i+1},x^{l(i+1)-1},x^{l(i+1)}\in C_{\rho}$.
+Then the descent lemma implies
+$f(x^{l(i+1)})-f(x^{i+1})\leq\frac{L_{\rho}}{2}\|x^{l(i+1)}-x^{i+1}\|^{2}+\left<\nabla f(x^{i+1}),x^{l(i+1)}-x^{i+1}\right>.$ (4.21)
+Meanwhile, from the fact that $\frac{1}{2}\|\cdot-x^{l(i+1)-1}\|^{2}$ is convex, we have
+$\frac{1}{2}\|x^{i+1}-x^{l(i+1)-1}\|^{2}-\frac{1}{2}\|x^{l(i+1)}-x^{l(i+1)-1}\|^{2}\leq\left<x^{i+1}-x^{l(i+1)-1},x^{i+1}-x^{l(i+1)}\right>$ (4.22)
+Putting (4.20), (4.21), and (4.22) together yields that
+$q(x^{l(i+1)})-q(x^{i+1})$ $\nabla f(x^{l(i+1)-1})-\nabla f(x^{i+1})+\gamma_{l(i+1)-1}\left(x^{i+1}-x^{l(i+1)-1}\right),x^{i+1}-x^{l(i+1)}\right>$ (4.23)
+$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ +\frac{L_{\rho}}{2}\|x^{l(i+1)}-x^{i+1}\|^{2}$
+$\leq\frac{L_{\rho}}{2}\|x^{l(i+1)}-x^{i+1}\|^{2}+\left(\bar{\gamma}_{\rho}+L_{\rho}\right)\|x^{i+1}-x^{l(i+1)-1}\|\|x^{i+1}-x^{l(i+1)}\|$
+$\leq\left(L_{\rho}+\frac{\bar{\gamma}_{\rho}}{2}\right)\|x^{l(i+1)}-x^{i+1}\|^{2}+\frac{\bar{\gamma}_{\rho}}{2}\|x^{l(i+1)}-x^{l(i+1)-1}\|^{2},$
+therefore, by the inequality $\sqrt{a+b}\leq\sqrt{a}+\sqrt{b}$ for all $a\geq 0$ and $b\geq 0$, as well as Proposition 4.5, one has
+$\sqrt{q(x^{l(i+1)})-q(x^{i+1})}$ $\leq\sqrt{L_{\rho}+\frac{\bar{\gamma}_{\rho}}{2}}\|x^{l(i+1)}-x^{i+1}\|+\sqrt{\frac{\bar{\gamma}_{\rho}}{2}}\|x^{l(i+1)}-x^{l(i+1)-1}\|$ (4.24)
+$\leq\left(c_{\hat{k}}\sqrt{L_{\rho}+\frac{\bar{\gamma}_{\rho}}{2}}+\sqrt{\frac{\bar{\gamma}_{\rho}}{2}}\right)\|x^{l(i+1)}-x^{l(i+1)-1}\|$
+for all $k_{0}\leq i\leq k+1$. Then, we have
+$\|x^{i+1}-x^{i}\|<\sqrt{\frac{2}{\mu}}\sqrt{q(x^{l(i+1)})-q(x^{i+1})}\leq\left(c_{\hat{k}}\sqrt{\frac{2L_{\rho}+\bar{\gamma}_{\rho}}{\mu}}+\sqrt{\frac{\bar{\gamma}_{\rho}}{\mu}}\right)\|x^{l(i+1)-1}-x^{l(i+1)}\|$
 for all $k_{0}\leq i\leq k+1$ and $i\in K$, which definitely says that
-$\|x^{i+1}-x^{i}\|\leq\sqrt{\frac{\bar{\gamma}_{\rho}+L_{\rho}}{\mu}}\|x^{l(i+1)-1}-x^{l(i+1)}\|\quad\forall k_{0}\leq i\leq k+1\ \text{and}\ i\in K.$
+$\|x^{i+1}-x^{i}\|\leq\sqrt{2}\left(c_{\hat{k}}+1\right)\sqrt{\frac{\bar{\gamma}_{\rho}+L_{\rho}}{\mu}}\|x^{l(i+1)-1}-x^{l(i+1)}\|\quad\forall k_{0}\leq i\leq k+1\ \text{and}\ i\in K.$
 We know that the induction hypothesis yields Theorem 4.8 holds automatically for $v=k+1$, hence, we have
-$\sum_{k_{0}=i\in K}^{k+1}\|x^{i+1}-x^{i}\|\leq\sqrt{\frac{\bar{\gamma}_{\rho}+L_{\rho}}{\mu}}\sum_{k_{0}=i\in K}^{k+1}\|x^{l(i+1)-1}-x^{l(i+1)}\|$ (4.22)
-$\leq\sqrt{\frac{\bar{\gamma}_{\rho}+L_{\rho}}{\mu}}\sum_{i=k_{0}}^{k+1}\|x^{l(i+1)-1}-x^{l(i+1)}\|$
-$<\frac{4m}{3}\sqrt{\frac{\bar{\gamma}_{\rho}+L_{\rho}}{\mu}}\left(\sqrt{\frac{2\big{(}q(x^{l(k_{0}-m-1)})-q(\bar{x})\big{)}}{\delta\gamma_{\min}}}+\frac{2(\bar{\gamma}_{\rho}+L_{\rho})}{\delta\gamma_{\min}}\chi\big{(}q(x^{l(k_{0})})-q(\bar{x})\big{)}\right).$
+$\sum_{k_{0}=i\in K}^{k+1}\|x^{i+1}-x^{i}\|\leq\sqrt{2}\left(c_{\hat{k}}+1\right)\sqrt{\frac{\bar{\gamma}_{\rho}+L_{\rho}}{\mu}}\sum_{k_{0}=i\in K}^{k+1}\|x^{l(i+1)-1}-x^{l(i+1)}\|$ (4.25)
+$\leq\sqrt{2}\left(c_{\hat{k}}+1\right)\sqrt{\frac{\bar{\gamma}_{\rho}+L_{\rho}}{\mu}}\sum_{i=k_{0}}^{k+1}\|x^{l(i+1)-1}-x^{l(i+1)}\|$
+$\leq\frac{4\sqrt{2}m\left(c_{\hat{k}}+1\right)}{3}\sqrt{\frac{\bar{\gamma}_{\rho}+L_{\rho}}{\mu}}\left(\sqrt{\frac{2\big{(}q(x^{l(k_{0}-m-1)})-q(\bar{x})\big{)}}{\delta\gamma_{\min}}}+\frac{2(\bar{\gamma}_{\rho}+L_{\rho})}{\delta\gamma_{\min}}\chi\big{(}q(x^{l(k_{0})})-q(\bar{x})\big{)}\right).$
 By (4.18) and (4.25), we have
 $\sum_{i=k_{0}}^{k+1}\|x^{i+1}-x^{i}\|\leq\sum_{k_{0}=i\in K}^{k+1}\|x^{i+1}-x^{i}\|+\sum_{k_{0}=i\in\overline{K}}^{k+1}\|x^{i+1}-x^{i}\|$
-$\leq\left(\frac{8m(\bar{\gamma}_{\rho}+L_{\rho})}{3\delta\gamma_{\min}}\sqrt{\frac{\bar{\gamma}_{\rho}+L_{\rho}}{\mu}}+\frac{m}{3}\sqrt{\frac{2(\bar{\gamma}_{\rho}+L_{\rho})}{\delta\gamma_{\min}-\mu}}\left(\frac{3}{m}+\frac{2(\bar{\gamma}_{\rho}+L_{\rho})}{\delta\gamma_{\min}}\right)\right)\chi\big{(}q(x^{l(k_{0})})-q(\bar{x})\big{)}$
-$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ +\left(\frac{4m}{3}\sqrt{\frac{\bar{\gamma}_{\rho}+L_{\rho}}{\mu}}+\frac{m}{3}\sqrt{\frac{2(\bar{\gamma}_{\rho}+L_{\rho})}{\delta\gamma_{\min}-\mu}}\right)\sqrt{\frac{2\big{(}q(x^{l(k_{0}-m-1)})-q(\bar{x})\big{)}}{\delta\gamma_{\min}}}$
-$\leq\left(\frac{8m(\bar{\gamma}_{\rho}+L_{\rho})}{3\delta\gamma_{\min}}\sqrt{\frac{\bar{\gamma}_{\rho}+L_{\rho}}{\mu}}+\frac{m}{3}\sqrt{\frac{2(\bar{\gamma}_{\rho}+L_{\rho})}{\delta\gamma_{\min}-\mu}}\left(\frac{3}{m}+\frac{2(\bar{\gamma}_{\rho}+L_{\rho})}{\delta\gamma_{\min}}\right)\right)\chi\big{(}q(x^{l(k_{0})})-q(\bar{x})\big{)}$
-$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ +\left(\frac{4m}{3}\sqrt{\frac{\bar{\gamma}_{\rho}+L_{\rho}}{\mu}}+\frac{m}{3}\sqrt{\frac{2(\bar{\gamma}_{\rho}+L_{\rho})}{\delta\gamma_{\min}-\mu}}+\sqrt{\frac{\delta\gamma_{\min}}{\delta\gamma_{\min}-\mu}}\right)\sqrt{\frac{2\big{(}q(x^{l(k_{0}-m-1)})-q(\bar{x})\big{)}}{\delta\gamma_{\min}}}$
+$\leq\left(\frac{8\sqrt{2}m\left(c_{\hat{k}}+1\right)(\bar{\gamma}_{\rho}+L_{\rho})}{3\delta\gamma_{\min}}\sqrt{\frac{\bar{\gamma}_{\rho}+L_{\rho}}{\mu}}+\frac{m}{3}\sqrt{\frac{2(\bar{\gamma}_{\rho}+L_{\rho})}{\delta\gamma_{\min}-\mu}}\left(\frac{3}{m}+\frac{2(\bar{\gamma}_{\rho}+L_{\rho})}{\delta\gamma_{\min}}\right)\right)$
+$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \cdot\chi\big{(}q(x^{l(k_{0})})-q(\bar{x})\big{)}$
+$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ +\left(\frac{4\sqrt{2}m\left(c_{\hat{k}}+1\right)}{3}\sqrt{\frac{\bar{\gamma}_{\rho}+L_{\rho}}{\mu}}+\frac{m}{3}\sqrt{\frac{2(\bar{\gamma}_{\rho}+L_{\rho})}{\delta\gamma_{\min}-\mu}}\right)\sqrt{\frac{2\big{(}q(x^{l(k_{0}-m-1)})-q(\bar{x})\big{)}}{\delta\gamma_{\min}}}$
+$\leq\left(\frac{8\sqrt{2}m\left(c_{\hat{k}}+1\right)(\bar{\gamma}_{\rho}+L_{\rho})}{3\delta\gamma_{\min}}\sqrt{\frac{\bar{\gamma}_{\rho}+L_{\rho}}{\mu}}+\frac{m}{3}\sqrt{\frac{2(\bar{\gamma}_{\rho}+L_{\rho})}{\delta\gamma_{\min}-\mu}}\left(\frac{3}{m}+\frac{2(\bar{\gamma}_{\rho}+L_{\rho})}{\delta\gamma_{\min}}\right)\right)$
+$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \cdot\chi\big{(}q(x^{l(k_{0})})-q(\bar{x})\big{)}$
+$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ +\left(\frac{4\sqrt{2}m\left(c_{\hat{k}}+1\right)}{3}\sqrt{\frac{\bar{\gamma}_{\rho}+L_{\rho}}{\mu}}+\frac{m}{3}\sqrt{\frac{2(\bar{\gamma}_{\rho}+L_{\rho})}{\delta\gamma_{\min}-\mu}}+\sqrt{\frac{\delta\gamma_{\min}}{\delta\gamma_{\min}-\mu}}\right)$
+$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \cdot\sqrt{\frac{2\big{(}q(x^{l(k_{0}-m-1)})-q(\bar{x})\big{)}}{\delta\gamma_{\min}}}$
 Hence, statement (b) holds for $k+1$, and this completes the induction.
@@
 Taking $k\to\infty$ in (4.14) therefore shows that $\{x^{k}\}$ is a Cauthy sequence and, thus, convergent.
-Since we already know that $\bar{x}$ is an accumulation point, it follows that the entire sequence $\{x^{k}\}$ is convergent to $\bar{x}$.
+Since we already know that $\bar{x}$ is an accumulation point, it follows that the entire sequence $\{x^{k}\}_{k\in\mathbb{N}}$ is convergent to $\bar{x}$.
 ∎
-Note that if $\overline{K}=\mathbb{N}$ or $m=0$, then we have $q(x^{k})>q(x^{k+1})$ for all $k\in\mathbb{N}$, i.e., $\{q(x^{k})\}$ is decreasing, then Algorithm 4.1 degenerates into [15, Algorithm 3.1], where the specific results about the convergence and convergence rate of the whole sequence were proposed.
-In the following, we illustrate the more general convergence rate of the whole sequence where $\overline{K}\neq\mathbb{N}$.
+Note that if $\overline{K}=\mathbb{N}$ or $m=0$, then we have $q(x^{k})>q(x^{k+1})$ for all $k\in\mathbb{N}$, i.e., $\{q(x^{k})\}_{k\in\mathbb{N}}$ is decreasing, then Algorithm 4.1 degenerates into [16, Algorithm 3.1], where the specific results about the convergence and convergence rate of the whole sequence were proposed.
+In the following, we illustrate the convergence rate of Algorithm 4.1 where $\overline{K}\neq\mathbb{N}$.
 Theorem 4.10.
-Let Assumption 3.2 and Assumption 4.2 hold, and let $\{x^{k}\}$ be any sequence generated by Algorithm 4.1.
-Suppose that $\{x^{k}\}_{\mathcal{K}}$ is a subsequence converging to some limit point $\bar{x}$, and that $q$ has the KL property at $\bar{x}$.
-Then the entire sequence $\{x^{k}\}$ converges to $\bar{x}$, and if the corresponding desingularization function has the form $\chi(t)=ct^{\theta}$ for some $c>0$ and $\theta\in(0,1]$, then the following statements hold:
+Let Assumption 3.2 and Assumption 4.2 hold, and let $\{x^{k}\}_{k\in\mathbb{N}}$ be any sequence generated by Algorithm 4.1.
+Suppose that $\{x^{k}\}_{k\in\mathcal{K}}$ is a subsequence converging to some limit point $\bar{x}$, and that $q$ has the KL property at $\bar{x}$.
+Then the entire sequence $\{x^{k}\}_{k\in\mathbb{N}}$ converges to $\bar{x}$, and if the corresponding desingularization function has the form $\chi(t)=ct^{\theta}$ for some $c>0$ and $\theta\in(0,1]$, then the following statements hold:
   1. (a)
-    if $\theta=1$, then the sequences $\{q(x^{l(k)})\}$ and $\{x^{k}\}$ converge in a finite numer of steps to $q(\bar{x})$ and $\bar{x}$, respectively.
+    if $\theta=1$, then the sequences $\{q(x^{l(k)})\}_{k\in\mathbb{N}}$ and $\{x^{k}\}_{k\in\mathbb{N}}$ converge in a finite numer of steps to $q(\bar{x})$ and $\bar{x}$, respectively.
   2. (b)
-    if $\theta\in[\frac{1}{2},1)$, then the sequence $\{q(x^{l(k)})\}$ converges Q-linearly to $q(\bar{x})$, and the sequence $\{x^{k}\}$ converges R-linearly to $\bar{x}$.
+    if $\theta\in[\frac{1}{2},1)$, then the sequence $\{q(x^{l(k)})\}_{k\in\mathbb{N}}$ converges Q-linearly to $q(\bar{x})$, and the sequence $\{x^{k}\}_{k\in\mathbb{N}}$ converges R-linearly to $\bar{x}$.
   3. (c)
@@
 $\big{(}q(x^{l(k+1)})-q(\bar{x})\big{)}^{2(1-\theta)}\leq\frac{1}{\xi}\big{(}q(x^{l(k)})-q(x^{l(k+1)})\big{)}\quad\forall k\geq k_{0}.$ (4.27)
-Since $\{q(x^{l(k)})\}$ is decreasing, then statements (a), (b), and (c) regarding the sequence $\{q(x^{l(k)})\}$ follow from [14, Lemma 2.14].
-It remains to verify the convergence rate with respect to the sequence $\{x^{k}\}$.
+Since $\{q(x^{l(k)})\}_{k\in\mathbb{N}}$ is decreasing, then statements (a), (b), and (c) regarding the sequence $\{q(x^{l(k)})\}_{k\in\mathbb{N}}$ follow from [15, Lemma 2.14].
+It remains to verify the convergence rate with respect to the sequence $\{x^{k}\}_{k\in\mathbb{N}}$.
 We now consider the different cases of $\theta$.
-  • $\theta=1$: for all $k\geq k_{0}$, (4.24) implies that $q(x^{l(k+1)})-q(x^{l(k)})\leq-\xi$, which yields that $\{q(x^{l(k)})\}$ converges to $q(\bar{x})$ in finite steps, then one has $\{x^{k}\}$ also converges to $\bar{x}$ in finite steps.
+  • $\theta=1$: for all $k\geq k_{0}$, (4.27) implies that $q(x^{l(k+1)})-q(x^{l(k)})\leq-\xi$, which yields that $\{q(x^{l(k)})\}_{k\in\mathbb{N}}$ converges to $q(\bar{x})$ in finite steps, then one has $\{x^{k}\}_{k\in\mathbb{N}}$ also converges to $\bar{x}$ in finite steps.
   • $\theta\in(0,1/2)$: for all $k\geq k_{0}$ and $s>k$, (4.14) implies that
@@
     $\leq(a+bc)\big{(}q(x^{l(k-m-1)})-q(\bar{x})\big{)}^{\frac{1}{2}}.$
-    Note that $\{q(x^{l(k)})\}$ Q-linearly converges to $q(\bar{x})$, then $\{x^{k}\}$ R-linearly converges to $\bar{x}$ (by taking $s\to\infty$).
+    Note that $\{q(x^{l(k)})\}_{k\in\mathbb{N}}$ Q-linearly converges to $q(\bar{x})$, then $\{x^{k}\}_{k\in\mathbb{N}}$ R-linearly converges to $\bar{x}$ (by taking $s\to\infty$).
 This completes the proof. ∎
 5 Conclusion
-This manuscript presented the convergence and rate-of-convergence results of two types of nonmonotine proximal gradient methods, i.e., combined with average line search and max line search, respectively.
-These results were established under the KL property, requiring only that the smooth part of the objective function has locally (rather than globally) Lipschitz continuous gradients.
-Note that the partitioning of the index set is a useful tool for the convergence of NPG.
-At least for NPG with average line search, the convergence theory is independent on the specific choice of partitioning.
-Additionally, our results in Section 4 address the question raised in [17, Final Remarks], demonstrating that the technique of proof there can indeed be extended to NPG with max line search.
+This manuscript presented the convergence and rate-of-convergence results of two types of NPG methods, i.e., those combined with average line search and max line search, respectively.
+The results were established under the KL property, requiring only that the smooth part of the objective function has locally (rather than globally) Lipschitz continuous gradients.
+Note that the partitioning of the index set is a useful tool for the convergence analysis of NPG methods.
+At least for the NPG method with average line search, the convergence theory is independent on the specific choice of partitioning.
+Additionally, our results in Section 4 addressed the question raised in [18, Final Remarks], demonstrating that the technique of proof there can be extended to the NPG method with max line search.